### v2

# The Vector Grounding Problem

Abstract
The remarkable performance of large language models (LLMs) on complex linguistic tasks has sparked debate about their capabilities. Unlike humans, these models learn language solely from textual data without directly interacting with the world. Yet they generate seemingly meaningful text on diverse topics. This achievement has renewed interest in the classical ‘Symbol Grounding Problem’ – the question of whether the internal representations and outputs of symbolic AI systems can possess intrinsic meaning that is not parasitic on external interpretation. Although modern LLMs compute over vectors rather than symbols, an analogous problem arises for these systems, which we call the Vector Grounding Problem. This paper has two main goals. First, we distinguish five main notions of grounding that are often conflated in the literature, and argue that only one of them, which we call referential grounding, is relevant to the Vector Grounding Problem. Second, drawing on philosophical theories of representational content, we provide two arguments for the claim that LLMs and related systems can achieve referential grounding: (1) through preference fine-tuning methods that explicitly establish world-involving functions, and (2) through pre-training alone, which in limited domains may select for internal states with world-involving content, as mechanistic interpretability research suggests. Through these pathways, LLMs can establish connections to the world sufficient for intrinsic meaning. One potentially surprising implication of our discussion is that that multimodality and embodiment are neither necessary nor sufficient to overcome the Grounding Problem.
1 Introduction
The remarkable performance of Large Language Models (LLMs) in sophisticated linguistic and cognitive tasks – once considered reliable indicators of human-like intelligence – has rekindled philosophical discussions on the nature of linguistic competence (Millière & Buckner 2024, Millière forthcoming). Despite being trained exclusively on text-based data without direct interaction with the physical world, these models exhibit the ability to generate coherent and contextually relevant paragraphs, answer questions, and solve commonsense reasoning tasks in a wide range of domains. This ability raises intriguing questions about the relationship between language, meaning, and reference in disembodied artificial systems that learn from input of raw text. [^1] For recent discussion, see Butlin (2021), Cappelen & Dever (2021), Pavlick (2022), Chalmers (2023), Mandelkern & Linzen (2024), Lederman & Mahowald (2024), Miracchi Titus (2024), Grindrod (2024), Borg (2025), Pepp (2025).
In this paper, we revisit the classical ‘Symbol Grounding Problem’, originally formulated for symbolic AI systems, in light of contemporary Generative Pre-Trained Models, and especially LLMs. The core issue can be summarized as follows: if AI systems are designed to process linguistic inputs without direct interaction with the world, can their internal representations and outputs possess any meaning beyond the interpretations that we, as intelligent beings embedded in the world, project onto them? We use the term intrinsic meaning to capture this notion of interpretation-independent meaning, distinguishing it from meaning that is imposed from outside the system itself, which is thus extrinsic. [^2] This follows Harnad (1990)’s terminology. Intrinsic meaning, as we understand it here, is fully compatible with externalist views about how states and outputs come to acquire interpretation-independent meaning and reference. The intrinsic/extrinsic distinction cuts across the internalism/externalism one. We thank an anonymous referee for prompting this clarification. We argue that a similar conundrum arises for Generative Pre-Trained Models, which we dub the Vector Grounding Problem. Since these systems compute over continuous vectors rather than discrete symbols, the problem assumes a slightly different form. After outlining this new version of the classical problem, we suggest that there are strong reasons to believe that a basic form of grounding is possible in current LLMs.
Specifically, we contend that the learning process of LLMs establishes functional roles for their internal states that go beyond mere linguistic prediction. Although LLMs learn from text-based data, they undergo selection pressures that can favour internal states tracking features of the world whenever extra-linguistic criteria influence the optimization objective. This can occur directly by fine-tuning LLMs on objectives that reflect extra-linguistic norms – such as factuality – or indirectly via the implicit extra-linguistic success conditions of latent tasks embedded within the training data or provided in context. When internal states that were selected to respond differentially to specific worldly conditions causally influence the generation of linguistic outputs, those outputs become grounded in the world in a way that makes them meaningful independently of human interpretation. The crucial question is thus not whether outputs can be interpreted as meaningful by humans (they obviously can), but whether the outputs themselves stand in the right kinds of relations to the world such that they would be meaningful even in the absence of external interpretation.
Importantly, this paper does not address whether LLMs and related models understand language (or anything at all), engage in linguistic acts, or possess cognitive or mental properties. [^3] See Goldstein & Levinstein (2024) and Chalmers (2025) for discussions that also draw on theories of content but consider whether LLMs can be ascribed propositional attitudes, and Stoljar & Zhang (2025) for a discussion of whether LLMs can think. We do not claim that these models ‘know’ what they mean, or even what meaning is, nor that they have any conception of truth or falsity. Our sole contention is that the problem of grounding internal states and outputs in the world is solvable in principle, and in some cases arguably already in practice, for LLMs and related models. While having grounded internal states may be a crucial step toward understanding, agency, knowledge, and various other cognitive and mental capacities, it is unlikely that they are, by themselves, sufficient for those capacities. This clarification is especially important, as the current debate occasionally blurs these distinctions – a conflation we strive to avoid in this paper.
The paper proceeds as follows. First, we introduce the Symbol Grounding Problem in the context of classical AI (section 2). Next, we present the Vector Grounding Problem as it applies to LLMs and related models (section 3). We then differentiate between various senses of grounding, showing that only one of them – referential grounding – is essential to the Vector Grounding Problem (section 4). Drawing on philosophical theories of representational content, we articulate a set of conditions that must be fulfilled by a system to achieve referential grounding (section 5). Finally, we argue that at least some LLMs and related multimodal models satisfy these requirements, and therefore have internal representations and outputs with intrinsic meaning (section 6).
2 The Classical Symbol Grounding Problem
Research on artificial intelligence has historically followed two competing paradigms: classical (or symbolic), and connectionist. The classical paradigm focuses on systems that represent and manipulate information using discrete symbols, based on explicit rules or procedures. Symbolic manipulation is purely syntactic; it is based on the shapes of symbols and the way in which they can be combined into complex symbolic structures (Simon & Newell 1971). In contrast, the connectionist paradigm focuses on artificial neural networks that process information in a distributed way through interconnected nodes, learning patterns from data rather than following explicit symbolic rules (Mcclelland et al. 1986).
The Symbol Grounding Problem, as identified by Harnad (1990), asks how the meaning of symbols in a formal symbol system can be intrinsic to the system itself, rather than merely parasitic on the meanings in the minds of external interpreters. More precisely:
Symbol Grounding Problem. How can symbol tokens, manipulated solely on the basis of their shapes according to syntactic rules, acquire meaning and refer to entities and properties in the real world – rather than merely relating other meaningless symbol tokens – without depending on the semantic interpretations provided by external meaning-bearing systems? [^4] It is common to take Harnad’s Symbol Grounding Problem as an elaboration of an earlier, highly influential criticism of AI: Searle (1980)’s Chinese Room thought experiment. Harnad (1990) himself followed this line.
Harnad (1990) provides a simple and powerful intuition pump to motivate the existence of a Symbol Grounding Problem in classical, symbolic AI: the Chinese Dictionary. Take a system, be it biological or artificial, which knows no natural language. Suppose that we want it to learn Chinese (or any other language), but the only resource made available to the system is a single-language dictionary, say a comprehensive Chinese-Chinese dictionary. Will the system be able to learn Chinese just by using the dictionary? The thought experiment invites the intuition that the system would not be able to do so. After all, the most it can do is to look up strings of symbols and see what other strings of symbols follow them. The former would be the words to be defined, the latter their definitions. However, by themselves the symbols are just meaningless shapes followed by other shapes, and it seems intuitively hopeless that the meaning of words (parts of those strings of symbols) could be learned this way. What seems to be missing is a way out from the symbolic ‘merry-go-round’: to connect those symbols in the dictionary to what they pick out in the world. But how can that be possible if the system has no access to the world, but only to further symbols?
Symbolic AI systems, according to Harnad, find themselves in this predicament. All they have access to are strings of symbols and how they follow each other. There seems to be no way for connections between such symbols and the world to be established, and thus no way for the symbols to have intrinsic meaning. Humans, who do have representations arguably acquired by means of engagement with the world, are, on the other hand, able to interpret those symbolic representations and outputs as meaningful. The meaning of the representations is hence extrinsic, depending on human interpreters. Take out the interpreter, and no meaning is to be found.
In brief, symbolic AI systems, the argument goes, manipulate symbols that are not connected to the world, that are not grounded in the world; and thereby have no intrinsic meaning. Classical AI systems, in other words, suffer from the Symbol Grounding Problem.
3 The Vector Grounding Problem
Connectionist models have come a long way since the 1980s, with deep neural networks (DNNs) emerging as powerful tools for various tasks (LeCun et al. 2015). Recent research has converged onto training deep neural networks on vast amounts of data to perform a broad range of tasks they were not explicitly designed for. Such DNNs are routinely called ‘Foundation Models’ (Bommasani et al. 2022), though we will use the more neutral label ‘Generative Pre-Trained Models’. Many among the most popular Generative Pre-Trained Models today are powered by the Transformer architecture, first introduced by Vaswani et al. (2017). Transformers use an attention mechanism to model global dependencies between sequential elements in the input, such as sentences in natural language. The general inductive biases supplied by this mechanism make Transformers suitable for many domains, including natural language processing and computer vision.
Some Generative Pre-Trained Models, called Large Language Models (LLMs), are trained exclusively on text. State-of-the-art LLMs like GPT-4  and Claude have over $10^{11}$ parameters and are trained on massive corpora spanning hundred of billions of words, containing a significant subset of all written text on the internet. After training, LLMs can generate well-written and coherent paragraphs about virtually any topic. Furthermore, given explicit instructions in natural language in the input sequence, or ‘prompts’, they can perform a broad range of tasks, including: summarising long-form content, translating between multiple languages, answering complex questions, engaging in commonsense reasoning, solving simple logic and maths problems, generating code for simple programs, or explaining jokes (Srivastava et al. 2023).
LLMs process input text into a sequence of tokens (i.e., full words, word-parts such as ‘-ing’, punctuation marks, and other linguistic symbols). LLMs encode each token as a vector of real-valued numbers – known as an embedding – in a high-dimensional vector space, and estimate the probabilities of all possible next tokens in the sequence based on the previous embeddings. The resulting prediction of the next token in the sequence is then compared to the actual next token, and the model’s parameters are subsequently updated to make the correct token more probable in that context, and other tokens less probable. While LLMs do not treat words as the main units of processing, the combination of tokens can form an indefinite number of possible words, and we can view the vectors for such composing tokens as jointly forming a word embedding. This training process (confusingly known as pre-training) encourages the model to produce context-sensitive embeddings, with semantically similar words having vectors that are close together in the high-dimensional space.
There is a deep connection between the way in which LLMs learn embeddings from linguistic corpora, and the old idea from structural linguistics according to which word meaning can be inferred from the contexts in which they appear, known as the ‘distributional hypothesis’ (Harris 1954, Firth 1957). In LLMs, semantic similarity is determined based on distribution across contexts: words that appear in similar contexts are modelled by nearby vectors. The dimensions of the vectors are tuned by the model to best predict surrounding words, and capture distributional patterns. Although individual dimensions are not typically interpretable, the positions of vectors in the space are a result of statistical patterns in the linguistic data, and reflect semantic similarities in light of how the words are used in context. [^5] Vision-Language Models (VLM) that can process images in addition to text work in a similar way: they encode text and images as vectors in a single high-dimensional space. We will discuss whether these models have specific implications for the grounding problem in subsection 7.3. In this way, embeddings in LLMs are not arbitrary or conventional – as the symbols in classical AI systems typically are – since they are shaped by statistical patterns in language.
However, LLMs appear to face a similar predicament to that of traditional symbolic AI systems when it comes to grounding and intrinsic meaning. Their reliance on continuous vectors, rather than discrete symbols, seems to make no difference to escaping the merry-go-round of representations, and being connected to the external world – LLMs are trapped in a carousel of numbers, rather than symbols. Much like their classical counterparts, LLMs manipulate vectors that do not seem grounded in the world and thus lack intrinsic meaning. Consequently, they grapple with an analogous difficulty, which we call the Vector Grounding Problem.
Mirroring Harnad’s 1990 Symbol Grounding Problem, the Vector Grounding Problem asks how the vector embeddings inside Generative Pre-Trained Models can have intrinsic meanings, rather than merely being parasitic on the meanings in the minds of external interpreters. More precisely:
Vector Grounding Problem. How can vector embeddings, manipulated solely based on their numerical values according to algebraic operations, acquire meaning and refer to entities and properties in the real world – rather than merely relating other meaningless vectors – without depending on the semantic interpretations provided by external meaning-bearing systems?
Both Grounding Problems, as stated here, focus on internal states, such as symbols and vector embeddings. If the Grounding Problems are solved, and such internal states come to have intrinsic meaning, we get straightforward reasons to think that the systems in question produce outputs that are intrinsically meaningful. For the meaning of the outputs would thus derive from the meaning of the system’s internal representations, rather than from the interpretations given by human users. [^6] In such a case, human users could misinterpret the system’s output if their interpretation does not conform to what the system actually represented – in a way akin to how humans can misinterpret each other.
It is worth noting that the Grounding Problems are particularly relevant to systems that generate novel outputs. This sets them apart from systems that merely reproduce outputs generated by humans, such as paper copiers. When a human-generated text is copied, there is no question about its intrinsic meaning: it was produced by a system with meaningful internal representations to convey a specific meaning to the reader. In contrast, generative systems like LLMs produce outputs that are not, for the most part, mere copies of human-generated text; they generate entirely novel sentences whose meanings cannot be equated to what any specific human has ever thought or said. Unlike with paper copiers, the question of whether LLM outputs are meaningful cannot be straightforwardly explained by appealing to a human author. [^7] We thank an anonymous referee for pressing us on this point.
Bender & Koller (2020) propose a thought-provoking thought experiment, the “Octopus Test”, to pump intuitions about the inescapability of the Vector Grounding Problem for LLMs:
The Octopus Test. Two English-speaking castaways find themselves stranded on neighbouring islands, separated by treacherous waters. Fortuitously, they discover telegraphs left by previous inhabitants, connected via an underwater cable, which enables them to communicate by exchanging telegraphic messages. Unbeknownst to them, a superintelligent octopus inhabits these waters and taps into the underwater cable, intercepting their messages. Though the octopus lacks any knowledge of English, its superintelligence allows it to detect statistical patterns in the telegraphic messages and to form an accurate representation of the statistical relationships between various telegraphic signals. The octopus decides to sever the underwater cable and position itself at the ends of the two resulting cable segments, receiving and replying to telegraphic signals from both castaways based on the statistical patterns it has identified. Whether or not the castaways notice this change, the messages sent by the octopus intuitively seem to hold no intrinsic meaning. After all, the octopus simply adheres to the statistical patterns it has learned from listening in on the previous exchanges between the humans, without any understanding of the human interpretation of the signals, such as ‘coconut’ or ‘tree’. Furthermore, the octopus likely does not comprehend that the signals possess meaning or serve a communicative function. [^8] Bender & Koller (2020) emphasise the fact that the octopus lacks communicative intent when generating telegraphic signals. Since our focus here is the narrower, more basic question of vector grounding, rather than the more complex capacities related to language comprehension and communication, we will set those considerations aside.
As should be clear, the octopus in this thought experiment is operating roughly as an extremely powerful LLM would. Regardless of how the octopus represents the signals and their statistical relations, be it through vectors or other means, it only has access to the signals and their statistical relations – without any connection to the entities and properties in the real world that the words encoded in those signals represent. Consequently, even if the octopus can produce outputs that appear sensible to humans in all situations, these outputs lack intrinsic meaning. [^9] Our version of the Octopus Test goes beyond Bender & Koller (2020) by assuming that it is possible the humans would never notice they were no longer communicating with each other. In contrast, Bender & Koller (2020) contend that the octopus would fail to produce appropriate outputs to questions involving concrete objects it has never perceived, such as ‘how do I build a coconut catapult?’. This claim seems to be challenged by the often plausible replies that current LLMs provide to this sort of question. We should expect the superintelligent octopus to generate even more reliably plausible responses. The Octopus Test implies that LLMs generate outputs as devoid of intrinsic meaning as those produced by the superintelligent octopus. [^10] Arguably, LLMs’ outputs would be even more devoid of intrinsic meaning than the octopus’, given that the latter is a biological organism with cognitive states, motivation, superintelligence, and other features that LLMs lack. Similar considerations apply to the comparison between LLMs and parrots (Bender et al. 2021). In the case of these animals, even if their outputs carry some meaning, they typically do not align with the meaning of the language they produce. It is humans, whether castaways or otherwise, who interpret these outputs as meaningful because they resemble meaningful words and sentences. LLMs have no access to coconuts, trees, or anything else beyond language. The vectors they receive and manipulate during internal processing are merely meaningless arrays of numbers, ungrounded in anything outside statistical patterns in language, just like the outputs they generate. In other words, the Octopus Test invites the intuition that all meaning in the outputs produced by LLMs is extrinsic, i.e., it lies in the eye of the beholder, their human interpreters. This would be so, we might add, because LLMs are susceptible to the Vector Grounding Problem.
Throughout the remainder of this paper, we will challenge this claim and counter the intuition primed by the Octopus Test. We aim to show that, contrary to first appearances, LLMs indeed possess the features required to generate meaningful outputs based on meaningful internal representations.
Before developing this argument further, it is essential to clarify the target notion of grounding.
4 Five notions of Grounding
Various notions of grounding have been invoked by cognitive scientists, philosophers, and AI researchers as central for endowing internal representations and outputs with meaning, occasionally leading to confusion. In this section, we endeavour to bring clarity to this debate by identifying five primary notions of grounding at play in discussions about intrinsic meaning in cognitive systems, both biological and artificial (Figure 1). Furthermore, we contend that only one of these five notions of grounding is crucial for addressing the Vector Grounding Problem. If LLMs possess features that enable them to satisfy the criteria for this particular type of grounding, they become candidates for having representations and outputs with intrinsic meaning. With this in mind, let us examine each notion of grounding in turn.
Figure 1: Five notions of grounding. A. Referential grounding: a lexical representation (<DOG>) is connected to its worldly referent (a dog). B. Sensorimotor grounding: a lexical representation (<DOG>) is connected to a sensory representation (image of a dog). C. Relational grounding: a lexical representation (<DOG>) is connected to other lexical representations (<PET>, <FURRY>). D. Communicative grounding: Two speakers calibrate their interpretation of an exchange to make use of the same lexical concept (<DOG>). E. Epistemic grounding: A lexical representation (<DOG>) is connected to information stored in a knowledge base (about dog facts).
4.1 Referential Grounding
The first notion of grounding, which we call referential grounding, closely aligns with the philosophical notion of reference. [^11] Partee (1981) uses the phrase “language-to-world grounding” to designate a closely related notion, which she contrasts with “language-to-language grounding” (see subsection 4.3 below for discussion of the latter). While Partee’s notion of language-to-world grounding focuses on what ties down the intensions of linguistic expressions with respect to their extra-linguistic content, our notion of referential grounding is broader, including the relation between internal representations and their worldly referents. It is the type of grounding that connects representations to things in the world. Reference can be loosely described as a relation that enables representations to ‘hook onto’ worldly entities or properties. [^12] Much philosophical work has been dedicated to examining the nature and varieties of reference. For our limited purposes, however, the intuitive notion captured by the ‘hooking’ metaphor will suffice. For example, the word ‘Paris’ refers to the city of Paris, and arguably, so does a map of Paris. Both representations allow us to receive and convey information about the actual city of Paris. Similarly, many of our thoughts seem to hook onto specific entities and properties in the world, allowing us to entertain thoughts that refer to Paris, such as beliefs about it being the capital of France or desires to visit the city.
Referential grounding, as we use the term, is whatever makes it so that representations, whether public (e.g., words and maps) or internal (e.g., beliefs and subpersonal states), hook onto the world. It is plausible that different kinds of representation require different means for hooking onto the world in the relevant sense, that is, to be referentially grounded in specific worldly entities and properties (Coelho Mollo 2015).
4.2 Sensorimotor Grounding
Another type of grounding that is often appealed to, especially in cognitive psychology, is what we call sensorimotor grounding. For an internal representation to be grounded in this sense is for it to be fundamentally linked to sensorimotor representations available within a cognitive system. For instance, the lexical representation ‘football’ is sensorimotorically grounded if its representation in a cognitive system can be reduced to, or is intrinsically connected to, sensorimotor representations of kicking movements, of the visual perception of a football pitch, of the auditory perception of the sound of a ball being kicked, etc. This kind of grounding plays a prominent role in theories in cognitive psychology that embrace a moderate form of embodied cognition, such as those positing the existence of “perceptual symbol systems” in human cognition (Barsalou 1999, Pulvermüller 1999).
Interestingly, sensorimotor grounding is also the solution proposed by Harnad (1990) to his Symbol Grounding Problem. He argued that the symbols in classical AI systems needed to be grounded in two types of representation: (a) iconic representations, which depict transformations of the projections of object ‘shapes’ on a system’s sensory surfaces; and (b) categorical representations, which represent invariant sensory features that facilitate the identification and discrimination of objects in different categories. Harnad hypothesised that connectionist approaches to AI would be better equipped to provide the necessary iconic and categorical representations, which could then complement symbolic AI systems and result in hybrid symbolic-connectionist systems.
We will come back to Harnad’s suggestion in subsection 4.6, arguing that it fails to solve the Symbol Grounding Problem, and does not help with the Vector Grounding Problem either.
4.3 Relational Grounding
A third notion of grounding, that we may call relational grounding, is occasionally used to refer specifically to intra-linguistic relationships, in accordance with the idea that a word’s meaning in a language is partly determined by its relations to other words. [^13] The label is somewhat under-descriptive, since all forms of grounding are relational, though in different ways. What we wish to underline, however, is that on this sense of grounding the meaning of representations depends on the relations they stand in to other representations in the same representational system. Partee (1981) introduces the notion of “language-to-language grounding”, referring to the way in which a language learner acquires knowledge about the meanings of words by learning semantic connections between words in the language. According to Partee, one key way in which this happens is through accepting certain sentences as true, particularly those that are regarded as impervious to empirical challenge within the linguistic community. For example, learning that “bachelors are unmarried males” helps ground the meaning of ‘bachelor’ through its relation to ‘unmarried’ and ‘male’.
Relational grounding is closely related to inferential role semantics (Block 1986, Brandom 1994). Inferential role semantics identifies the meaning of a linguistic expression with its inferential relations to other expressions. For example, the word ‘and’ plays a logical role in reasoning that determines when compound statements are true or false based on the truth or falsity of their components. Likewise, words like ‘larger’ or ‘slower’ have inferential roles in comparisons, allowing us to determine comparisons between items or the relative properties of a single item. Knowledge of such relations between words underlies our ability to draw material inferences in language and thought.
Finally, there is a clear connection between Partee’s “language-to-language grounding” and distributional semantics. Distributional semantics seeks to analyse word meanings in terms of word-to-word relationships. However, instead of drawing on linguistic analyses of word meaning, it relies on analysing co-occurrence statistics in text corpora. Consequently, distributional semantic models capture relationships between words grounded in patterns of use, rather than linguistic theory. The high dimensionality of vector spaces in LLMs makes it possible to represent complex, multifaceted similarities between words – some of which plausibly reflect conceptual structure and/or inferential roles (Piantadosi & Hill 2022).
4.4 Communicative Grounding
The notion of grounding is also used in relation to another aspect of language, which has to do with communication (Clark & Brennan 1991, Traum 1994). In this context, the ‘grounding problem’ refers to the problem of establishing common ground between speakers in conversation. From this perspective, linguistic communication can be seen as a form of coordinated action that involves collaborating to reach a common understanding of what is said. The process of communicative grounding may involve explicit clarification strategies, such as prompting an interlocutor to repeat or reformulate a garbled or ambiguous statement. It also involves tracking communicative intentions during the exchange, to infer what is communicated from what is said. Such inferences can be made partly on the basis of pragmatic presuppositions, namely, what speakers take for granted or assume to be true when they use certain sentences in the context of the conversation (Stalnaker 2002).
The communicative notion of grounding has also been discussed by natural language processing (NLP) researchers (Brennan 1998, Di Maro 2021, Chandu et al. 2021). For example, Chandu et al. (2021) distinguish between static grounding, in which common ground in basic communication between a human user and a virtual agent is automatically established by querying a database (e.g., asking for a weather report); and dynamic grounding, in which common ground is more gradually established in back-and-forth communication between a human and a dialogue agent. Bender & Koller (2020) define meaning to be a relation between linguistic expressions and communicative intent, and characterise this relation as a form of grounding.
While communicative and referential grounding are distinct notions, they are related in important ways. Referential grounding anchors language to the world, but the same linguistic expressions may be referentially grounded in different ways for different speakers, due to differences in perception, knowledge, and conceptualisation. This means that calibrating reference in conversation requires communicative grounding – language users must work together to negotiate a common way of connecting language to the world. Without such collaboration, there is no guarantee that speakers really mean the same thing, even if they are using the same words.
4.5 Epistemic Grounding
Finally, the notion of grounding is occasionally used by NLP researchers to refer more minimally to the relationship between linguistic expressions and information stored in knowledge bases (Poon 2013, Tsai & Roth 2016, Chandu et al. 2021, Thoppilan et al. 2022). For example, an expert system may contain a knowledge base of facts and rules belonging to a particular domain. When the system receives user input, it can query this knowledge base to determine an appropriate response. A classic example is MYCIN, an expert system for diagnosing infectious blood diseases (Shortliffe 1977). When given patient symptoms, MYCIN would search its knowledge base of medical rules and facts to suggest possible diagnoses and recommend treatments. In this system, language is grounded in the structured knowledge contained in the expert system’s knowledge base. We call this notion epistemic grounding.
There are two ways in which current LLMs may involve epistemic grounding. Firstly, they can be hooked up to external databases used to retrieve information (Borgeaud et al. 2022, Thoppilan et al. 2022). Secondly, it has been suggested that even LLMs that are not augmented with retrieval capabilities might be functionally similar to knowledge bases themselves (Petroni et al. 2019). Indeed, they may encode a significant amount of world knowledge in their weights after pre-training, as illustrated by the remarkable performance of state-of-the-art LLMs on a wide range of factual question-answering benchmarks (OpenAI 2023, Bubeck et al. 2023).
4.6 What Grounding for LLMs?
We have identified five distinct notions of grounding that play prominent roles in discussions about meaning grounding in biological and artificial systems. Which of these notions is necessary for LLMs to possess representations and produce outputs endowed with intrinsic meaning? (Recall that intrinsic meaning is meaning that is independent from external interpretation.) To identify the most suitable notion of grounding for this task, we will employ an elimination strategy, examining each notion for its adequacy, beginning with Harnad’s1990 proposal of appealing to sensorimotor grounding as a solution to the Symbol Grounding Problem.
Sensorimotor grounding connects conceptual or linguistic representations to sensorimotor representations within the same cognitive system. This sort of grounding can be useful to shed light on some features of the former representations, how they work, and what roles they can play in the economy of a cognitive system. However, sensorimotor grounding falls short in addressing the merry-go-round problem, as it merely grounds one type of representation in another: it does not tackle the question of how sensorimotor representations establish a connection with the world. One might argue that sensorimotor representations are grounded in the world by resembling, to some extent, the entities they represent. However, the success of such an approach relies on an additional factor beyond sensorimotor grounding itself: it is the relevant resemblance relations that are supposed to ground those sensorimotor representations in the world. As sensorimotor grounding does not provide a connection with the world but only with other representations, it does not address the Symbol Grounding Problem. Similarly, this form of grounding does not resolve the Vector Grounding Problem, as grounding vectors in other vectors still leaves open the question of how any of those vectors come to be connected to the world.
The same line of reasoning applies to relational, communicative, and epistemic grounding. Relational grounding anchors linguistic representations in their relationships to other linguistic representations, suggesting that a representation’s meaning is determined by its connections to the meanings of other representations within a network of inferential, conceptual, or statistical relations. Although this kind of grounding clearly does not escape the representational carousel, inferential-role semantics may suffice for connecting at least some words to what they are about, e.g., logical connectives, as their meanings are arguably exhausted by the inferences they are involved in (Block 1986, Fodor 1990). Logical connectives are however a rather special case, as they are not supposed to represent entities in the world, but rather abstract relations between any possible entities. [^14] Therefore, if an LLM can learn the structure of relations that constitute the meaning of logical connectives and linguistic conjunctions from its training data, then its representations of these terms might be as intrinsically meaningful as they can be. Inferential-role semantics can thereby account, at best, for only a very small and special subset of the representations and outputs that biological and artificial systems possess and produce.
Communicative grounding seeks to anchor representations in the shared content of linguistic interactions between agents. However, this shared content relies on the representations within the agents’ cognitive systems and the meanings of any public representations used in the exchange. As with sensorimotor grounding, one type of representation is grounded in another, perpetuating the representational carousel without establishing a connection to the world beyond the representations.
Epistemic grounding, on the other hand, aims to ground linguistic representations in knowledge bases. Yet, this approach encounters a familiar obstacle: the representations within the knowledge bases themselves call for an explanation of how they acquire intrinsic meaning. Consequently, epistemic grounding alone fails to provide a comprehensive solution to the Grounding Problem, as it does not address the question of how representations come to possess intrinsic meaning to start with.
In sum, the form of grounding that underlies the other four types and is most relevant to addressing the Grounding Problems is what we have called referential grounding. Referential grounding anchors a representation – be it a symbol, a map, or a vector – in the world itself, rather than in other representations. It stands as the sole form of grounding that escapes the representational merry-go-round and connects a representation to whatever worldly entity or property it is about. [^15] Or almost: see the caveat about inferential-role semantics and logical connectives above. Consequently, to determine whether Generative Pre-Trained Models such as LLMs manipulate and produce representations and outputs with intrinsic meaning, we must assess whether they can achieve referential grounding. To do so, we first need to examine the requirements for a representation to be referentially grounded.
5 Theories of Representational Content: Causes and History
Philosophers of cognitive science have dedicated considerable effort over the past several decades to developing theories that explain how mental and cognitive states, such as beliefs, desires, and neural activations, can be about, or mean, entities and properties in the world. What mental and cognitive processes are about, or mean, is their ‘representational content’, or ‘content’ for short. The philosophical project of developing theories of representational content has the aim of illuminating how meaning can emerge in a world that otherwise seems devoid of it. After all, natural objects such as rocks and lakes generally do not possess internal states endowed with meaning and, by themselves at least, do not mean anything.
Some internal representations are basic, in the sense that their meaning, or content, does not derive from other states already endowed with meaning. They can be seen as the building blocks of more complex representations, which draw some of their meaning from those basic representations. For instance, ‘edge detectors’ in primary visual cortex (V1) are good candidates for being basic internal representations, whereas ‘shape detectors’ in downstream visual areas are not, since their contents derive partly from the contents of the basic edge detectors. Similarly, words and sentences are not basic in this sense, since their meaning derives from other, more basic representational states, such as beliefs, desires and intentions – themselves complex representations.
The distinction between basic and complex representations cuts across the distinction between intrinsic and extrinsic meaning. As a reminder, states with intrinsic meaning are those that carry meaning independently of an interpreter with mental states, such as a human, attributing meaning to them. Both the representations in primary visual cortex and the words and sentences humans utter are cases of representation with intrinsic meaning, though the former are basic while the latter are complex. By contrast, words randomly formed by the movement of ants in the sand, outputs from a random sentence generator – and, according to Bender & Koller (2020), the outputs of LLMs – only have extrinsic meaning: they are merely shapes that require human interpretation to convey any meaning at all.
We cannot here do justice to the rich and nuanced interdisciplinary research that has been carried out for the past several decades on the notion of cognitive representation. [^16] See Neander (2017), Millikan (2017), Shea (2018) for recent book-length defences of the main contenders today. What matters for our limited purposes is that, despite some disagreements on the finer details, philosophers of cognitive science have largely reached a consensus on the key factors that enable internal states in a system to become representations, both basic and complex (Shea 2018, Coelho Mollo 2022).
The first, most fundamental point of agreement is that an account of basic representational content must identify one or more natural relations between internal states and external entities and properties responsible for making the former be about the latter. These natural relations must not be ones that already involve states with meaning. Since the goal is to account for the emergence of meaning in the world, theories of basic content must exclusively rely on meaning-less states and processes to avoid vicious circularity. Much of the work in this area has been dedicated to identifying those relations that can yield fruitful content ascriptions and contribute to explaining the behaviour of cognitive systems.
Two natural relations are widely considered to be central to endowing basic internal states with meaning: causal-informational relations, and historical relations. In general terms, causal-informational relations occur when one state of the world conveys information about another due to a causally-generated correlation between the two states. A classic example is the relationship between smoke and fire: smoke is correlated with fire because it is caused by it, so the presence of smoke carries information about the presence of fire. Neurons in brain area V1 convey information about edges in the perceived visual scene, as their activation correlates (through a relatively complex causal chain) with the presence of edges in the world in front of the perceiving subject.
Causal-informational relations capture the main rationale for why systems possess basic internal representations: they allow systems to be sensitive to specific features of their environments – those for which internal states stand in causal-informational relations to. However, the mere existence of a causal-informational relation is insufficient to provide an account of basic representational content. An additional requirement often mentioned is that the causal-informational relation be exploited, that is, used by the system due to its information-carrying nature (Shea 2018). In other words, a system must use the presence of smoke to obtain information about the presence of fire, form a belief about there being a fire, guide fire-related behaviour, and so on, for smoke to be a viable candidate for representing fire for that system.
Similarly, the information about the world that edge detectors in V1 carry into the mammalian brain must be used downstream for further processing – such as detecting food sources – so as to guide appropriate behaviour. To be a representation necessarily involves playing a certain role: that of conveying information about the world and being used by the system in virtue of that. This further requirement complements the explanation for why internal representations are useful to systems: they are used by the system as sources of information about the world. It also helps to avoid the risk of trivialising the notion. Causal-informational relations are abundant in nature: every effect carries information about its cause. If every such relation should suffice for being a representation, representations would be so extremely trivial that their special explanatory power for cognition would be lost.
Exploited causal-informational relations alone, however, are still not enough to base an adequate theory of representational content, for another crucial feature of internal representations is missing: their capacity for truth or falsehood, correctness or incorrectness. To be a representation inherently involves the possibility of misrepresenting – the possibility of getting things wrong and failing to fit to the world accurately. V1 neurons may become active when no edge is present, as in the famous Kanizsa’s triangle illusion, thus conveying information to the system that, if not corrected by other subsystems, can lead to inappropriate behaviour (such as trying to touch the non-existent edge). Representations involve a kind of descriptive normativity (Neander 2017). That is to say, representations can represent successfully or unsuccessfully, better or worse. This descriptive normativity extends to external representations, such as sentences, and meaning more generally. They are kinds of things that can be true or false, accurate or inaccurate, appropriate or inappropriate. The rationale for systems to possess representations involves that they be accurate enough, enough of the time. Otherwise, representations would lead systems astray and hinder appropriate behaviour, rather than the other way around. Generally, representations have the function to represent correctly and accurately (to a degree of required precision). [^17] In some special cases, systematic misrepresentation can be a sensible price to pay. If the stakes are high enough, it makes sense for representations to err on the side caution: if you are trying to detect predators, it makes sense to produce many false-positive (misrepresenting predators as present when there are none), rather than even one, potentially fatal false-negative.
As the notion of representation is central for cognitive (and computer) science, the notion of function is central to biology. This has led philosophers of biology to develop scientifically-grounded, objective theories of function Wright (1973), Millikan (1989), Neander (1991), Godfrey-Smith (1994), Garson (2019). If functions exist out there in the world, they must be a matter of natural features and relations, rather than human interpretation. The dominant view of functions is the selected-effects theory Millikan (1989), Neander (1991), Garson (2019). On this view, functions are ensembles of the historical causes that explain the persistence of a certain trait or feature in a system. For instance, hearts today have the function to pump blood because past hearts, by pumping blood, causally contributed to the survival and reproduction of the systems having them. Pumping blood is what explains, causal-historically, why hearts still exist – and it is thereby their function. The relevant causes can be at the level of the species, typically through natural selection; or at the level of the individual, through selection-like learning processes during development.
The most influential views on representation incorporate the selected-effects theory to account for internal states having the function to carry information about the world and to guide appropriate behaviour Neander (2017), Millikan (2017), Shea (2018), Coelho Mollo (2022). The guiding idea is that, through selection and/or learning, states and processes in cognitive systems come to acquire functions they are supposed to perform. For example, neurons in V1 have the function of detecting edges in the visual scene because they were selected for this purpose during evolution via natural selection; and such selection occurred due to the causal-informational relations they maintain with actual edges in the world. When these neurons successfully detect edges, they fulfil their function, representing correctly; when they fail to fire in the presence of an edge or fire in the absence of one, they malfunction, misrepresenting. In brief, internal representations are states that carry information about the world, have the function to do so, and are appropriately used by the system in virtue of the information they carry.
Existing theories of basic representation vary in the emphasis they place on causal-informational or historical relations, as well as in the specific types of such relations they prioritise. Additionally, they may diverge on the inclusion of further relations, such as structural resemblance relations Shea (2018). Nevertheless, all of these theories, in one way or another, assert that a combination of these relations is central to the emergence of meaning in the world. To possess intrinsic meaning, at a minimum, is to stand in a specific set of causal-informational and historical relations with the thing in the world being represented. Theories of content, therefore, provide the essential ingredients that determine referential grounding, enabling internal states to hook onto things in the world. [^18] Philosophical discussions about cognitive representation have traditionally given considerable weight to metaphysical worries regarding the determinacy of content. Once seen instead as scientific, explanatory theories, we take such worries to be misguided (cf., Shea 2018, Coelho Mollo 2022).
6 Grounding Generative Pre-Trained Models
Philosophical theories of representation identify sets of relations that imbue representations with content, giving them intrinsic meaning. These primarily include causal-informational and historical relations. Our aim is to show that Generative Pre-Trained Models, such as LLMs, can have internal states and generate outputs with intrinsic meaning, i.e., that are referentially grounded in the world. [^19] For recent discussions that apply theories of representational content to LLMs, see Harding (2023), Grzankowski (2024), Goldstein & Levinstein (2024), Chalmers (2025). To begin, we will examine whether appropriate causal-informational relations exist between LLM representations/outputs and entities and properties in the world.
6.1 Causal-informational relations
Although LLMs are trained exclusively on linguistic data, thus precluding any direct causal-informational relationship with extra-linguistic entities, the linguistic data itself is generated by humans who stand in causal-informational relations to worldly entities and properties. Those data are shaped by humans’ interactions with the world. Among its various functions, language enables humans to express facts about the world, as well as their experiences, emotions, and ideas concerning it. The extensive corpus of text upon which LLMs are trained bears the ‘imprint’ that the world has left on language and linguistic expression.
A strict separation of form and meaning has been a central assumption in much of linguistics, from Ferdinand de Saussure’s principle of the “arbitrariness of the linguistic sign” (de Saussure 1916) to contemporary formal theories of syntax and semantics (Chomsky 1957, Montague 1970). However, this notion has faced growing criticism since the 1980s, especially in light of empirical investigation of the iconicity of language (Perniss & Vigliocco 2014). In this context, iconicity refers to a non-arbitrary relationship between the form of a linguistic sign and its referent, based on some similarity or analogy (Haiman 1985). The nature of this similarity relation can vary: imagistic iconicity involves a perceptual resemblance between a word and its referent (e.g., in onomatopoeias), while diagrammatic iconicity involves a structural resemblance .
Diagrammatic iconicity is pervasive in linguistic structures. For example, the ordering of events in narrative sequences tends to reflect closeness and succession in time (e.g. “Veni, vidi, vici”); what is nearest to the speaker in a literal or metaphorical sense tends to be mentioned first (e.g., “here and there”); elements whose referents are closely related in some way or other tend to be adjacent in sentences (principle of adjacency); the use of subordination reflects causality or conditionality between states-of-affairs (e.g., “If you study hard, you will pass the exam”), and so on (Van Langendonck 2010). The pervasiveness of diagrammatic iconicity in linguistic structures suggests that the way in which linguistic signs are organised and structured is not arbitrary, and reflect some ways in which the world is organised and structured.
While diagrammatic iconicity is not typically characterised in distributional terms, patterns of co-occurrence in language reflect stable statistical regularities in the way in which linguistic elements are used together. Certain words tend to occur together more frequently than others, and certain grammatical constructions tend to co-occur with particular semantic categories. These patterns of co-occurrence reflect the underlying structure of the world that speakers are trying to describe. Thus, contra Bender & Koller (2020), form and meaning are far from being so easily separable, especially at the level of general patterns of language use.
The patterns of language use derived from the training corpus reveal important aspects of the world. They embed information about patterns of interaction between humans and the world, human ways of categorising their environments, and the selection of features that are more or less salient to them. Indeed, a plausible hypothesis for LLMs’ success in a wide range of tasks is that they have induced general semantic patterns from their training data, which closely resemble what could be termed semantic ‘laws’. [^20] For some preliminary work on the potential form of such semantic laws, see Wolfram (2023). These laws are not merely formal or syntactic; rather, they tap into the intricate relationships between linguistic form and the world it describes. [^21] See Merrill et al. (2024) for concrete evidence that LLMs can induce entailment relationships between sentences from co-occurrence statistics.
In short, LLMs are trained on data that is shaped by humans and their interactions with the world. LLMs stand in no direct causal connection to worldly entities, but they do stand in indirect ones, mediated by the humans who generated the training data. Importantly, while the complete lack of direct causal relations to the world sets LLMs apart from humans, reliance on indirect causal relations does not. Many human representations are also acquired by indirect, mediated ways. Indeed, testimony and social learning underlie a considerable amount of the internal representations humans come to learn, and are considered to be key processes that help explain cultural knowledge and cultural evolution – in their turn central factors for the species’ success Tomasello (2009), Sterelny (2014), Heyes (2018).
These include representations of entities one has not had any direct causal contact with. For example, most people have never been in direct causal contact with enriched uranium, but that does not entail that their thoughts and statements about uranium lack referential grounding – they are indeed about uranium. Representations of culturally constructed notions, such as ‘democracy’ or ‘university’, are also acquired this way – they rely on complex sets of social practices and conventions that are transmitted not by direct causal contact with documents or buildings, but rather by social learning.
In a way akin to many human representations, the internal representations in LLMs are indirectly connected to the world through testimony and ‘social’ learning. Since their training data contains innumerable instances of human communication, LLMs should be seen as an additional link in the chain of social information transmission that endows representations with causal-informational relations to worldly entities, and thus partly underpins intrinsic meaning.
In a sense, thereby, the causal-informational relationships between LLMs and the world are dependent on humans. This dependence is however distinct from the one that drives the Vector Grounding Problem. The Problem is not how the representations of connectionist models come to acquire intrinsic meaning, but whether these representations have intrinsic meaning at all – instead of being merely ascribed meaning by external interpreters. To take once again the example of enriched uranium, LLMs find themselves in a position similar to most humans: their representations and outputs about enriched uranium intrinsically refer to enriched uranium – irrespective of any interpreter’s gloss – even though the process enabling such reference depends on a chain of transmission that involves other systems with intrinsic meaning. In our terminology, such internal states are non-basic representations with intrinsic meaning.
In sum, LLMs possess the necessary elements to fulfil the causal-informational requirement for referential grounding. They stand in complex, mediated causal-informational relations to the world. These connections are mediated since they pass through human linguistic production, where language primarily functions as a tool to express aspects of the world. [^22] This analysis would extend, mutatis mutandis, to multimodal inputs, which are also mediated by human interactions with the world (e.g., photographs are captured by humans with a camera). However, one might argue that some worldly properties instantiated by visual samples in a multimodal dataset, such as shapes, can be accessed more directly than they could through the mediation of language. This is because images (and samples in other sensory domains) are entirely iconic, while language only exhibits limited patterns of iconicity. We discuss whether multimodal models fulfil the historical requirement for representational content in subsection 7.3 below. They are complex due to the intricate causal-informational relationships between the patterns of language use, the semantic ‘laws’ that govern them, the structure of the world they have (culturally) evolved to capture, and the causal chains of communication and testimony that produce the learning data – of which the LLM itself can be considered a link. In light of these considerations, we contend that Generative Pre-Trained Models, such as LLMs, can meet the causal-informational criteria necessary for possessing referentially grounded representations and outputs.
6.2 Historical Relations
Although it is important to differentiate between causal-informational and historical relations when considering referential grounding, these aspects are often intertwined in real-world cases. The causal-informational relations previously discussed with respect to LLMs possess a historical dimension, as they depend on patterns of interaction between humans and their environments. Likewise, in many of the interesting cases pertaining to human cognition, the meaning-endowing causal-informational relationships can be traced back to past episodes of learning. That said, as discussed in section 5, finding causal-informational relations between LLMs and the world is the easiest part of our investigation, given how abundant such relations are. Much more challenging is to show that LLMs have internal states that not only carry information about worldly entities, but that also have the function to carry such information – so that the LLM can use it in generating appropriate outputs.
Selection processes and learning are the historical mechanisms that theories of representation commonly invoke to explain how representations come to have the relevant function. These processes explain how a system acquires a specific function or purpose, in this case, the production and/or use of representations. Such functions clarify how cognitive systems can represent accurately – when they produce and use representations that fulfil the relevant function – as well as misrepresent, when the representations produced and/or used are inappropriate for the task. While it is widely accepted that Generative Pre-Trained Models undergo a learning-like process during training, it is less clear whether they undergo the relevant kind of selection process. It can be argued that they do, at least to the same degree as other human artefacts (Artiga 2016, Coelho Mollo 2019), given that numerous design and implementation choices form an integral part of creating artefacts such as AI systems.
The crux of the matter, however, lies in determining whether LLMs stand in historical relations to the right sorts of things, namely the entities and properties in the world that could ground their representations Butlin (2021). At first glance, the situation appears analogous to that of causal-informational relations: LLMs stand in historical relations to the world, albeit in a mediated manner, through the human-generated data used for training them. The critical question is: can the internal states and outputs of LLMs acquire the function to represent extra-linguistic entities, despite their being trained only on language? In other words, the question is whether LLMs can acquire the sort of world-involving functions needed for referential grounding, and which underlie the descriptive normativity that is an essential feature of meaning.
To address this question, it is important to distinguish between the proximate and ultimate functions performed by LLMs during training and inference. The proximate function of LLMs is plainly next-token prediction: predicting the next token in a text sequence, given the context provided by that sequence. This proximate function is infra-linguistic, and so are the standards for its fulfilment or failure, namely, whether the LLM predicts the correct tokens reliably enough. While such function may be suitable for underpinning relational grounding, it appears inadequate for contributing to the world-involving normativity necessary for referential grounding. In contrast, ultimate functions refer to the broader, more abstract goals that the model is tasked with achieving through next-token prediction. These functions can be world-involving in principle, such as generating text that accurately describes real-world states of affairs, or solving problems based on real-world information. In what follows, we will consider several potential sources for world-involving ultimate functions in the training and use of LLMs.
Our general argument goes as follows. The training process of LLMs is a form of selection – through learning – that stabilises the functional roles of the model’s internal states. This process selects for internal states whose ultimate functional role extends beyond mere linguistic prediction to include extra-linguistic, world-involving functions. This is most evident in models that undergo fine-tuning to satisfy human preferences, since such fine-tuning explicitly involves selecting internal states that increase the probability of outputs satisfying world-involving norms, such as factual accuracy. However, we will also present evidence suggesting that pre-training alone can, in some contexts, select for internal states with world-involving content, albeit in a more indirect way.
If LLMs indeed have internal states with extra-linguistic intrinsic meaning, and if those states causally influence the generation of linguistic outputs, then those outputs are thereby referentially grounded in the world. This would mean that at least some outputs of LLMs are intrinsically meaningful in the sense discussed earlier: their meaning does not depend on the interpretations projected onto them by human users.
The key premise in this argument – that the training process of LLMs can select for internal states with extra-linguistic functional roles – requires detailed defence. In what follows, we offer two complementary arguments supporting this premise. The first, which we call the argument from post-training, focuses on how preference tuning methods explicitly establish world-involving functional roles for LLM internal states. The second, which we call the argument from pre-training, examines how even the basic next-token prediction learning objective might, under certain conditions, implicitly select for ultimate world-involving functions.
6.2.1 The Argument from Post-Training
The training process of modern LLMs, particularly those deployed in consumer-facing chatbots like ChatGPT or Claude, involves two distinct phases: pre-training and post-training. Pre-training involves training the base model from scratch on a next-token prediction objective across vast corpora of text. By contrast, post-training adapts this base model to behave in ways that go beyond generating merely plausible text completions, making LLMs more useful for various downstream tasks and applications. The post-training phase typically includes two types of fine-tuning: instruction tuning, and preference tuning.
Instruction tuning is a form of supervised fine-tuning on curated instruction-response pairs (Ouyang et al. 2022). Given an instruction, the model is trained to generate the desired output by minimising the difference between its predictions and the target responses in the instruction dataset. This teaches the model to generate appropriate responses that follow user instructions rather than simply generating plausible completions of text sequences.
Preference tuning, by contrast, trains LLMs to generate outputs that better align with selected normative criteria, such as helpfulness, harmlessness, and honesty or factual accuracy (Askell et al. 2021). Unlike standard supervised fine-tuning, which only demonstrates what to do, preference tuning explicitly trains models to distinguish between better and worse responses according to these criteria. This is typically implemented by training on pairs of responses where human evaluators have indicated a preference for one over the other.
Reinforcement Learning from Human Feedback (RLHF) is the most well-known preference tuning method (Christiano et al. 2017). RLHF begins by collecting human preferences over model outputs for a diverse set of prompts. These preferences are used to train a reward model that predicts how humans would evaluate various responses. The LLM is then fine-tuned using reinforcement learning to maximize this reward function while remaining reasonably close to its original behaviour. Another increasingly popular approach is Direct Preference Optimization (DPO), which achieves similar results without explicitly modelling the reward function (Rafailov et al. 2024). Instead, DPO directly optimises the model to assign higher probabilities to preferred responses and lower probabilities to rejected ones.
When an LLM is optimised according to human preferences that incorporate extra-linguistic normative standards like factual accuracy, this creates a selection process that reinforces specific activation patterns based on their relationship to worldly states. These internal states acquire a specific functional role: to track relevant aspects of the world that enable the model to generate outputs that satisfy human preferences for truthfulness.
To illustrate how preference tuning establishes world-involving functions, consider a toy example focused solely on factual accuracy using DPO. We begin with a language model that has undergone both pre-training and instruction tuning. This model can follow instructions reasonably well, but has not been specifically optimised for factual accuracy. To align the model with the norm of factual accuracy, we collect a dataset of human preferences where judges rank responses based solely on which one contains more truthful information and fewer falsehoods.
When we apply DPO to this model, it adjusts the model’s parameters to increase the probability of generating factually accurate responses and decrease the probability of generating less accurate ones. [^23] Note that DPO also incorporates a constraint that prevents the model from deviating too far from its original behaviour, to maintain the right balance between preference satisfaction and fluency. Through this process, the model’s internal states are selected specifically because they enable the model to satisfy the norm of factual accuracy, thereby providing a standard of success that depends not on linguistic patterns alone, but on the correspondence between the generated text and states of affairs in the world.
This functional role provides another piece of the story – in addition to causal-informational relations – of how LLM internal states can be referentially grounded, and thus overcome the Vector Grounding Problem. The internal states of the preference-tuned LLM represent aspects of the world because they were selected specifically for their ability to carry information about those aspects in service of generating truthful outputs. When the model succeeds in generating a factually accurate response, its internal representations are fulfilling their function of tracking relevant worldly states.
Importantly, the accuracy of the model is not what determines whether its internal states are referentially grounded. Even a partially inaccurate model prone to occasional hallucinations can have internal states with world-involving content, provided those states were selected for their ability to carry information about aspects of the world. What matters for grounding is not perfect accuracy, but rather the presence of a representational function with world-involving correctness conditions. Preference tuning thus provides a clear mechanism by which LLMs can acquire internal states with extra-linguistic functional roles, insofar as RLHF or DPO incorporate factual accuracy as an explicit criterion over rewarded preferences. [^24] At least in those domains where such factual accuracy has been emphasized during preference tuning.
While we focus on factual accuracy, all three common norms in preference tuning – the “3 Hs” of helpfulness, harmlessness, and honesty – are arguably world-involving. Helpfulness requires generating outputs that genuinely assist users with their goals, which depends on real-world task success conditions rather than mere linguistic coherence. Harmlessness involves avoiding outputs that could lead to real-world negative consequences, requiring sensitivity to causal relationships between linguistic outputs and potential worldly outcomes. Even instruction tuning alone could potentially establish world-involving functions, as successfully following instructions often requires tracking aspects of the world that would make the instruction completion successful. Many instructions implicitly contain success conditions that depend on extra-linguistic factors. Be that as it may, factual accuracy provides the most direct and straightforward case of world-involving content, where the connection between internal states and worldly conditions is the most evident. It offers the clearest case for how LLM internal states can be referentially grounded, and thus endowed with intrinsic meaning.
6.2.2 The Argument from Pre-Training
We now turn to a more controversial claim: that pre-training alone can, in certain contexts, establish world-involving functions.
During pre-training, LLMs are explicitly trained to minimise the negative log-likelihood of the next token given previous tokens – an objective limited to the statistical distribution of tokens in the training corpus. [^25] See McCoy et al. (2024) for an approach to studying LLMs, which they call the “the teleological approach”, that takes as its starting point this proximate function. They show that LLM behaviour is strongly shaped by the distributional statistics of the training dataset. As we argue here, however, this is compatible with LLMs also, at least in some cases, fulfilling ultimate, world-involving functions. Sensitivity to distributional patterns can be a means to fulfilling ultimate goals, given all the worldly information embedded in such patterns. At first glance, this appears to be a purely infra-linguistic task with no connection to the world. However, as we have seen in subsection 6.1, the patterns of language use in the training corpus are not arbitrary; they reflect systematic regularities in how humans use language to communicate about the world.
From an information-theoretic perspective, there are vastly more ways to linguistically express world states than there are world states themselves. For next-token prediction, two strategies are possible: direct memorisation of observed sequences, or latent variable factorisation that learns to represent underlying causal factors. The parameter efficiency of the latter strategy relative to the former scales with the ratio of linguistic expressions to world states – typically a very large number. This creates a strong selection pressure for internal states that track latent variables, as they enable more efficient prediction across diverse linguistic contexts.
The Information Bottleneck principle provides a formal framework for understanding this process (Tishby & Zaslavsky 2015). Neural networks learn to extract approximate, minimally sufficient statistics of the input with respect to the output. This involves finding a maximally compressed mapping of the input that preserves as much information as possible about the output. In language models, this connection is even more fundamental. As shown by Delétang et al. (2024), minimising the next-token prediction loss is mathematically equivalent to finding the most efficient way to compress tokenised information. For complex data distributions generated by structured world domains, the most efficient compression involves factorising the problem according to the causal structure that generated the data.
Research in mechanistic interpretability provides concrete evidence that models pre-trained on next-token prediction can develop internal representations with extra-linguistic content at least in some domains. One particularly interesting case is Othello-GPT, a transformer model trained solely to predict moves in Othello game transcripts (Li et al. 2023). Despite receiving no explicit information about Othello’s rules or board structure, this model was found to develop internal representations that encode the current game state in a player-centric manner. Nanda et al. (2023) showed through intervention studies that specific directions in the model’s activation space encode board positions as "Mine," "Yours," or "Empty." When they intervened on these representations by ’flipping’ a tile from ’yours’ to ’mine’ in the model’s activation space, the model’s predictions changed in ways consistent with the altered board state. This suggests not only that the model encodes the board state, but that this encoding plays a causal role in determining its predictions of legal next moves. [^26] For a longer philosophical discussion of causal interventions used to understand interrnal states of Othello-GPT, see Millière & Buckner (forthcoming).
The Othello-GPT example illustrates how selection processes during pre-training can establish ultimate functional roles with extra-linguistic correctness conditions. During training, the model developed specific directions in its activation space that linearly encode the board state from the perspective of the current player. These linear representations emerged because they enabled the model to accurately predict legal moves across diverse game contexts. Their function became to carry information about the current state of the Othello board, with accuracy conditions dependent on the actual game state rather than merely on linguistic patterns. This example is particularly compelling because game transcripts are systematically constrained by the rules of Othello – to predict legal moves effectively, the model developed a linearly decodable internal model that tracks the underlying game state.
The case of LLMs trained on diverse corpora is undoubtedly more complex than Othello-GPT. Games represent a special class of domains with respect to the grounding problem. Unlike open-ended conversation, where success criteria may be fuzzy and context-dependent, games typically have explicit, formalised rules that define success and failure categorically. These rules create a clearly bounded domain with a fixed ontology (game pieces, positions, legal moves), and deterministic success conditions. By contrast, general LLMs encounter messy, open-ended data with few rigid underlying rules. To the extent that language models may acquire internal states with world-involving content, similarly to Othello-GPT with respect to the game world, this presumably only happens in limited domains and contexts, in which next-token prediction on text sequences provides the right kind of learning signal. That said, there is converging evidence from mechanistic interpretability research that LLMs have internal representations that carry information about discourse entities and their relations, and that such representations are causally relevant to the generation of outputs (e.g., Li et al. 2021; see Millière & Buckner forthcoming for a review).
A useful perspective for understanding how pre-training can establish world-involving functional roles is to view it through the lens of meta-learning. Meta-learning – often characterized as ‘learning to learn’ – is the process by which a model improves its learning capabilities through experience across multiple tasks, essentially learning how to learn more efficiently. The pre-training of LLMs can be viewed as a form of meta-learning because it trains the model on a diverse distribution of tasks implicitly contained within the next-token prediction objective. When predicting the next token in various contexts across a large corpus, the model encounters many implicit ‘tasks’, ranging from syntactic completion to factual recall, reasoning, and various domain-specific problems. This defines a bi-level optimisation structure characteristic of meta-learning: an outer loop (the training process itself) that optimises model parameters, and an inner loop (the forward pass during inference) that applies these learned parameters to solve new tasks on the fly (Wu & Varshney 2024). This helps explain why pre-trained LLMs can rapidly adapt to new tasks from just a few examples at inference time, a capacity known as in-context learning (ICL): they have essentially learned optimisation strategies that generalise across the diverse task distribution encountered during pre-training, enabling efficient transfer to novel contexts during inference.
Many different approaches to understanding in-context learning have been proposed. One prominent theory suggests that ICL involves implicitly optimising over latent task objectives. For example, von Oswald et al. (2024) showed that autoregressive transformers trained on sequence prediction tasks develop gradient-based optimisation algorithms in their forward pass. This suggests that when pre-trained LLMs encounter a series of examples in their context window, they can perform an implicit optimisation process over the latent task space. In other words, they can identify the underlying task structure from the examples, and construct an internal objective that captures the task function. When the in-context examples reflect an extra-linguistic task, the implicit objective function defined by the task may also be extra-linguistic. [^27] See, for example, Patel & Pavlick (2022)’s experiment which involves prompting the model in a few-short learning regime to learn mappings from colour terms like ‘cyan’ to points in RGB colour space such as [0, 100, 100]. This is an instance of a world-involving function, having a standard of success and failure directly contingent on how the world is. Optimising this implicit function during the forward pass should dynamically select activation patterns that increase the likelihood of outputs satisfying the world-involving task constraints.
In summary, the argument from pre-training suggests that even without post-training methods like preference tuning, language models may develop internal representations with ultimate world-involving functions through the selective process of next-token prediction over structured text data, at least in some limited domains. This can occur because inducing the latent causal structure of the data provides an informationally efficient solution to the prediction task across diverse contexts. Evidence from mechanistic interpretability studies supports this view at least in toy domains, showing that Generative Pre-Trained Models can develop internal states that carry information about extra-linguistic properties and causally influence outputs. Furthermore, the meta-learning perspective on in-context learning suggests that pre-trained LLMs can implicitly optimise ultimate extra-linguistic task functions at inference time without parameter update.
7 Implications
7.1 Parameter-identical duplicates
The historical component of our account of referential grounding invites consideration of a classic philosophical objection. The ‘swampman’ thought experiment envisions a perfect duplicate of a human spontaneously formed by random chance (in the original thought experiment, by a lightning strike on a swamp), and thus without any causal history. We can construct an analogous thought experiment for LLMs:
Swamp LLM. Suppose a language model $M_{1}$ is trained normally, undergoing both pre-training on next-token prediction and post-training with instruction tuning and preference tuning. Now imagine that an untrained language model $M_{2}$ is randomly initialised and, by extraordinary coincidence, happens to have exactly the same parameter values as the fully trained $M_{1}$, despite never having been exposed to any training data.
If $M_{1}$ achieves referential grounding on our view, does $M_{2}$ also achieve it? Our answer aligns with the standard response to the swampman scenario: despite being parameter-identical to $M_{1}$, $M_{2}$ would lack referential grounding. First, the internal states of the swamp system, at least at the moment of creation, would lack any causal-informational relations to the world, given that, by hypothesis, the system has no causal history. Moreover, the absence of a selection and learning history means that the internal states of the swamp system lack any functions, including functions to carry information about the world Shea (2018). Consequently, there is no principled way to determine which outcomes count as successes versus failures for a swamp system. $M_{2}$ might produce outputs that appear meaningful to human interpreters, but without the selective history that shaped $M_{1}$, there is no objective basis for treating some of its outputs as ‘correct’ and others as ‘incorrect’. The parameter values themselves, divorced from the process that selected them, cannot ground representational content. [^28] This scenario is somewhat akin to the example, mentioned above, of ants forming the shape of words in the sand by pure chance; or of chimps eventually reproducing a Shakespeare play by randomly hitting keys in a typewriter. Such outputs have no intrinsic meaning, as they were not produced in the appropriate ways, though referentially grounded beings, such as ourselves, may extrinsically impose meaning on them.
It is worth noting, however, that referential grounding could be established relatively quickly if $M_{2}$ were subsequently fine-tuned. Even a brief learning history would begin to establish the relevant world-involving functions Shea (2018). What matters for grounding is not the parameter values per se, but the historical process through which those parameters were selected to perform specific functions.
We can also consider a variant of this thought experiment: [^29] We thank an anonymous referee for raising this scenario.
Lucky Pre-Training. Suppose $M_{1}$ is trained normally, undergoing both pre-training and post-training. $M_{2}$ undergoes only pre-training on next-token prediction, but through some extraordinary coincidence, ends up with exactly the same parameter values as the fully trained $M_{1}$.
This scenario differs importantly from the standard swampman case, as $M_{2}$ does have a learning history – it just differs from that of $M_{1}$. The answer to whether $M_{2}$ achieves referential grounding depends on whether one accepts our argument from pre-training in Section 6.2.2.
If one accepts that pre-training alone can, in some cases, establish world-involving functional roles (as we argued), then $M_{2}$ could achieve referential grounding despite skipping the post-training phase. However, the fact that $M_{2}$ happens to have identical parameter values to $M_{1}$ would be a mere coincidence irrelevant to the grounding question. Indeed, its referential grounding would be limited to that achieved via pre-training, while the internal states that drew their grounding from preference tuning in $M_{1}$ would be ungrounded in $M_{2}$.
These thought experiments highlight the crucial role that historical relations play in referential grounding. The internal states of a system, no matter how complex or functionally organised, cannot have representational content without the appropriate history. At the same time, our position avoids an overly restrictive view of what learning histories can establish referential grounding, allowing that different training regimes might enable a system to overcome the Vector Grounding Problem through different causal-historical paths.
7.2 The impact of fine-tuning
In a recent paper, Grindrod (2024) raises an objection to our argument from post-training. Grindrod acknowledges our claim that fine-tuning with human feedback introduces world-involving functions that could in principle establish the necessary historical relations for referential grounding. However, he objects:
The problem with this idea is that it would be odd to claim that a relatively marginal procedure that is designed to optimize the performance of a LLM on specific tasks ends up being the difference between intentionality and its lack. The reason, simply put, is that the fine-tuning procedures – whether RLHF, supervised training, or something else – are not the reason why we have such impressive LLMs today. (Grindrod 2024, p. 71)
Grindrod’s objection rests on the premise that it would be implausible for a “relatively marginal procedure” to be the critical factor determining whether LLMs possess intrinsic meaning. He points out that pre-trained models without preference tuning already demonstrate impressive capabilities, suggesting that whatever makes LLMs impressive cannot be the same factor that provides them with intrinsic meaning.
This objection misunderstands the purpose of our appeal to preference tuning. We are not claiming that preference tuning explains the impressive capabilities of LLMs or their ability to produce contextually appropriate outputs. Rather, we are specifically addressing how LLMs can overcome the Vector Grounding Problem – how their internal states can possess intrinsic meaning by being referentially grounded in the world. These are distinct questions.
Intrinsic meaning does not determine the complexity or impressiveness of outputs, neither in biological nor in artificial systems. Many relatively simple biological systems possess internal states with intrinsic meaning, while many complex systems might generate impressive outputs without intrinsic meaning. The quality of outputs and their status as intrinsically meaningful are orthogonal issues.
What matters for referential grounding is not how much a particular training procedure contributes to performance quality, but whether it establishes the right kind of functional relations between internal states and worldly entities. Even if preference tuning procedures make only marginal contributions to an LLM’s overall capabilities, they can fundamentally alter the normative conditions under which the model operates by introducing world-involving correctness standards.
7.3 Multimodality
The preceding discussion has identified two potential sources of referential grounding for LLMs: (1) preference tuning with human feedback that incorporates world-involving criteria such as factual accuracy, and (2) pre-training on next-token prediction, which may select for internal states with world-involving functional roles in limited domains where this is the most efficient way to achieve good prediction accuracy.
When we turn to multimodal and embodied systems, we need to examine whether and how these sources of referential grounding extend to such systems. Our analysis leads to a surprising conclusion: multimodality and embodiment are neither necessary nor sufficient conditions for referential grounding in artificial systems, though they can complement and potentially enhance referential grounding when the right kind of learning process is available.
Let us first consider a simple multimodal model like CLIP (Radford et al. 2021). CLIP is developed by jointly training an image encoder and a text encoder to predict the correct pairings of images and text captions using a contrastive learning objective. During training, CLIP receives batches of image-text pairs and learns to maximize the similarity between embeddings of matched pairs while minimizing the similarity between embeddings of unmatched pairs. This training approach allows CLIP to develop representations that capture semantic relationships between visual and linguistic content. Through contrastive pre-training with image-caption pairs, CLIP achieves what we called sensorimotor grounding, albeit to a very limited extent compared to biological organisms – namely connecting linguistic vector representations to pictorial vector representations. CLIP also arguably achieves some degree of relational grounding in the purely linguistic domain, although there is evidence that contrastive learning from image-caption pairs severely limits the degree to which this kind of model can induce syntactic structure compared to text-only LLMs (Yuksekgonul et al. 2023).
As far as causal-informational relations go, the account we gave for LLMs should also apply to models like CLIP: since both images and their captions are shaped by human interactions with the world, CLIP should stand in complex, mediated causal-informational relations to the world. The images in CLIP’s training data were produced by humans using cameras to capture real-world scenes, and the captions were written by humans to describe those scenes. This establishes an indirect causal chain connecting CLIP’s internal representations to worldly entities and properties. However, the contrastive learning objective optimises purely for matching images with their captions, without establishing clear world-involving functional roles for CLIP’s internal states. There is no obvious selection pressure for internal states that track features of the world beyond what is necessary for the matching task, and no clear normative standard for success beyond the matching task itself. Thus, whether a model like CLIP can achieve referential grounding through the pathway outlined in the argument from pre-training is doubtful.
By contrast with CLIP, some multimodal models are built with a pre-trained language model backbone. For example, typical VLMs that can receive images as well as text as input, such as Flamingo and LLaVA, leverage existing LLMs (with frozen weights) and learn a mapping between the visual encoder and the LLM (Bordes et al. 2024). This approach is computationally more efficient than training both vision and text encoders from scratch, as it only requires learning the connection between modalities. To the extent that the LLM backbone achieves referential grounding from pre-training on text alone, our framework entails that such model should inherit this property. The addition of a superficial mapping to the embedding space of a visual encoder provides a form of sensorimotor grounding, but it does not contribute to the referential grounding inherited from the pre-trained LLM.
7.4 Embodiment
Embodiment is often mentioned as the best way to secure that AI systems are grounded in the world, via their robotic bodies. Some embodied systems like SayCan (Ahn et al. 2022) integrate a pre-trained LLM with a robotic body in a way that keeps these components functionally separate. In SayCan, an LLM breaks down verbal instructions into sequences of ’atomic’ behaviours that the robot can enact through low-level visuomotor control. Importantly, the LLM component is not further trained or fine-tuned based on the success or failure of the robot’s actions. In such systems, the embodied aspect adds nothing to the referential grounding of the language model component. Without additional training that would establish new causal-historical relations between the LLM’s internal states and the outcomes of physical actions, the LLM’s capacity for referential grounding remains exactly what it was before integration with the robotic system. Like multimodality, physical embodiment alone does not automatically contribute to referential grounding. What matters is the causal-historical relationship established through training that selects for internal states with world-involving functional roles. Without such a learning history, embodiment remains external to the representational system itself.
When fine-tuning is involved, the story is a bit different. Recent advances in robotic learning have led to the development of sophisticated multimodal systems that integrate vision, language, and action. Models such as RT-2 (Zitkovich et al. 2023), RT-X (O’Neill et al. 2024), and OpenVLA (Kim et al. 2024) typically build upon language models or vision-language models that were pre-trained on text or multimodal data before being fine-tuned for robotic control. These models incorporate a pre-trained language model component and may inherit referential grounding from it. For instance, RT-2 builds upon vision-language models like PaLI-X and PaLM-E, which themselves incorporate language models pre-trained on text data. To the extent that these underlying language models achieve referential grounding through the mechanisms we have identified, the system as a whole may inherit this grounding.
However, VLA models go beyond merely inheriting referential grounding from their language model components. During embodied training, they are specifically trained to map visual inputs and natural language instructions to robotic actions that achieve specific goals in the physical world. This additional training process plausibly establishes new functional roles for the model’s internal states that are directly tied to successful physical interactions.
From our perspective, embodied training potentially enhances referential grounding by establishing or reinforcing causal-historical relations between internal states and features of the world that are relevant for successful manipulation. The training process selects for internal states that carry information about these features precisely because doing so improves performance on the physical tasks. This introduces an additional source of normativity tied to success in physical interaction, complementing whatever referential grounding was inherited from the pre-trained components.
8 Concluding Remarks
We have argued that the Symbol Grounding Problem, as introduced by Harnad (1990), can be adapted to apply to deep neural networks such as Generative Pre-Trained Models in we have called the Vector Grounding Problem. Despite initial intuitions, we have sought to demonstrate that at least a subset of current Generative Pre-Trained Models have the features necessary to be endowed with representations and outputs grounded in the world.
While our discussion has focused on LLMs and, to a lesser extent, on their multimodal counterparts, any AI system fulfilling the causal-informational and historical relations requirements delineated above can possess representations with intrinsic meaning. Meeting these requirements is far from trivial, but nonetheless within reach of current AI models. As philosophical theories of representation have shown, in order to understand how cognitive systems – both biological and artificial – function, we must resist the temptation to assume that the presence of intrinsic meaning necessarily entails complex cognitive abilities such as understanding, knowledge, or consciousness. While having states imbued with intrinsic meaning might constitute one of the fundamental conditions for possessing these other cognitive abilities, it is insufficient on its own. Activations of specific populations of neurons in the mammalian cortex may represent the presence of edges in the visual scene, but these neuronal activations do not understand or possess knowledge about edges. We suggest that a similar claim can be made regarding the internal representations and outputs of LLMs and related connectionist models.
We argued for this claim in two steps. First, we distinguished between various forms of grounding, which are often conflated in the cognitive science and computer science literature. We suggested that one particular form of grounding, referential grounding, is central to both the Vector Grounding Problem and the older Symbol Grounding Problem. Referential grounding pertains to the way in which representations hook onto the world. It is the most fundamental type of grounding, and arguably serves as the basis for all other notions of grounding (with the partial exception of relational grounding). Referential grounding requires two main factors: causal-informational relations between internal states and the world, and historical relations that imbue those states with functions to carry information about the world.
Second, we have shown that certain LLMs trained exclusively on linguistic data, particularly those that undergo preference tuning, are referentially grounded. These models benefit from indirect causal-informational relations with worldly entities and properties, mediated by human interactions with the world, which in turn shape the linguistic inputs they receive during training and inference. Preference tuning introduces an additional learning objective that is directly world-involving, as it encompasses a requirement for truthfulness or factuality. The normativity of this requirement supplies an extra-linguistic accuracy standard for the representations and outputs of LLMs. More tentatively, we have suggested that even without preference tuning, the representations and outputs of LLMs may achieve referential grounding in some limited domains, through pre-training and in-context learning on specific kinds of input sequences, such as those describing game moves.
This discussion led us to a counter-intuitive conclusion: some multimodal models trained on images in addition to text, like CLIP, are poor candidates for referential grounding, since they lack the kind of learning process that could supply world-involving functions. In fact, even some embodied systems that combine a ‘frozen’ LLM with a robotic body, like SayCan, also fail to acquire internal states and outputs that are referentially grounded. One benefit of our analysis is to highlight that multimodality and embodiment are neither necessary nor sufficient conditions for referential grounding in artificial systems, although they are certainly compatible with it if the right kind of learning process is available.

### v3

@@
 Abstract
-The remarkable performance of large language models (LLMs) on complex linguistic tasks has sparked debate about their capabilities.
-Unlike humans, these models learn language solely from textual data without directly interacting with the world.
-Yet they generate seemingly meaningful text on diverse topics.
-This achievement has renewed interest in the classical ‘Symbol Grounding Problem’ – the question of whether the internal representations and outputs of symbolic AI systems can possess intrinsic meaning that is not parasitic on external interpretation.
-Although modern LLMs compute over vectors rather than symbols, an analogous problem arises for these systems, which we call the Vector Grounding Problem.
-This paper has two main goals.
-First, we distinguish five main notions of grounding that are often conflated in the literature, and argue that only one of them, which we call referential grounding, is relevant to the Vector Grounding Problem.
-Second, drawing on philosophical theories of representational content, we provide two arguments for the claim that LLMs and related systems can achieve referential grounding: (1) through preference fine-tuning methods that explicitly establish world-involving functions, and (2) through pre-training alone, which in limited domains may select for internal states with world-involving content, as mechanistic interpretability research suggests.
-Through these pathways, LLMs can establish connections to the world sufficient for intrinsic meaning.
-One potentially surprising implication of our discussion is that that multimodality and embodiment are neither necessary nor sufficient to overcome the Grounding Problem.
+Large language models (LLMs) produce seemingly meaningful outputs, yet they are trained on text alone without direct interaction with the world.
+This leads to a modern variant of the classical symbol grounding problem in AI: can LLMs’ internal states and outputs be about extra-linguistic reality, independently of the meaning human interpreters project onto them?
+We argue that they can.
+We first distinguish referential grounding—the connection between a representation and its worldly referent—from other forms of grounding and argue it is the only kind essential to solving the problem.
+We contend that referential grounding is achieved when a system’s internal states satisfy two conditions derived from teleosemantic theories of representation: (1) they stand in appropriate causal-informational relations to the world, and (2) they have a history of selection that has endowed them with the function of carrying this information.
+We argue that LLMs can meet both conditions, even without multimodality or embodiment.
 1 Introduction
-The remarkable performance of Large Language Models (LLMs) in sophisticated linguistic and cognitive tasks – once considered reliable indicators of human-like intelligence – has rekindled philosophical discussions on the nature of linguistic competence (Millière & Buckner 2024, Millière forthcoming).
-Despite being trained exclusively on text-based data without direct interaction with the physical world, these models exhibit the ability to generate coherent and contextually relevant paragraphs, answer questions, and solve commonsense reasoning tasks in a wide range of domains.
-This ability raises intriguing questions about the relationship between language, meaning, and reference in disembodied artificial systems that learn from input of raw text.
-[^1] For recent discussion, see Butlin (2021), Cappelen & Dever (2021), Pavlick (2022), Chalmers (2023), Mandelkern & Linzen (2024), Lederman & Mahowald (2024), Miracchi Titus (2024), Grindrod (2024), Borg (2025), Pepp (2025).
-In this paper, we revisit the classical ‘Symbol Grounding Problem’, originally formulated for symbolic AI systems, in light of contemporary Generative Pre-Trained Models, and especially LLMs.
-The core issue can be summarized as follows: if AI systems are designed to process linguistic inputs without direct interaction with the world, can their internal representations and outputs possess any meaning beyond the interpretations that we, as intelligent beings embedded in the world, project onto them?
+The remarkable performance of Large Language Models (LLMs) in sophisticated linguistic and cognitive tasks—once considered reliable indicators of human-like intelligence—has rekindled philosophical discussions on the nature of linguistic competence.
+[^1] See Milliere2024 for a general introduction to language models, and Milliereforthcoming for a discussion of their relevance to debates about linguistic competence Despite being trained exclusively on text data without direct interaction with the physical world, these models exhibit the ability to generate coherent and contextually relevant paragraphs, answer questions, and solve commonsense reasoning tasks in a wide range of domains.
+This ability raises intriguing questions about the relationship between language, meaning, and reference.
+[^2] For recent discussion, see Butlin2021, cappelenMakingAIIntelligible2021, Pavlick_2022, chalmersDoesThoughtRequire2023, mandelkernLanguageModelsWords2024, Lederman2024, miracchititusDoesChatGPTHave2024, grindrodLargeLanguageModels2024, borgLLMsTuringTests, peppReferenceIntentionsLarge.
+In this paper, we revisit the classical ‘Symbol Grounding Problem’ (Harnad1990), originally formulated for symbolic AI systems, in light of contemporary LLMs.
+The core issue can be summarized as follows: if AI systems process linguistic inputs without direct interaction with the world, can their internal representations and outputs possess any meaning beyond the interpretations that we, as intelligent beings embedded in the world, project onto them?
 We use the term intrinsic meaning to capture this notion of interpretation-independent meaning, distinguishing it from meaning that is imposed from outside the system itself, which is thus extrinsic.
-[^2] This follows Harnad (1990)’s terminology.
+[^3] This follows Harnad1990’s terminology.
 Intrinsic meaning, as we understand it here, is fully compatible with externalist views about how states and outputs come to acquire interpretation-independent meaning and reference.
 The intrinsic/extrinsic distinction cuts across the internalism/externalism one.
-We thank an anonymous referee for prompting this clarification.
-We argue that a similar conundrum arises for Generative Pre-Trained Models, which we dub the Vector Grounding Problem.
-Since these systems compute over continuous vectors rather than discrete symbols, the problem assumes a slightly different form.
+We argue that a similar conundrum arises for LLMs, which we call the Vector Grounding Problem.
+Since these systems compute over continuous vectors rather than discrete symbols, the problem takes a somewhat different form.
 After outlining this new version of the classical problem, we suggest that there are strong reasons to believe that a basic form of grounding is possible in current LLMs.
-Specifically, we contend that the learning process of LLMs establishes functional roles for their internal states that go beyond mere linguistic prediction.
-Although LLMs learn from text-based data, they undergo selection pressures that can favour internal states tracking features of the world whenever extra-linguistic criteria influence the optimization objective.
-This can occur directly by fine-tuning LLMs on objectives that reflect extra-linguistic norms – such as factuality – or indirectly via the implicit extra-linguistic success conditions of latent tasks embedded within the training data or provided in context.
-When internal states that were selected to respond differentially to specific worldly conditions causally influence the generation of linguistic outputs, those outputs become grounded in the world in a way that makes them meaningful independently of human interpretation.
+Specifically, we contend that the learning process of LLMs endows some of their internal states with functions that go beyond mere linguistic prediction.
+Although LLMs learn from text data, they undergo selection processes that favour internal states tracking features of the world whenever extra-linguistic criteria influence the training objective.
+This can occur directly by fine-tuning LLMs on objectives that reflect extra-linguistic norms—such as factuality—or indirectly via the implicit extra-linguistic success conditions of latent tasks embedded within the training data or provided in context.
+When internal states selected to carry information about specific worldly conditions causally influence the generation of linguistic outputs, those outputs become grounded in the world in a way that makes them meaningful independently of human interpretation.
 The crucial question is thus not whether outputs can be interpreted as meaningful by humans (they obviously can), but whether the outputs themselves stand in the right kinds of relations to the world such that they would be meaningful even in the absence of external interpretation.
 Importantly, this paper does not address whether LLMs and related models understand language (or anything at all), engage in linguistic acts, or possess cognitive or mental properties.
-[^3] See Goldstein & Levinstein (2024) and Chalmers (2025) for discussions that also draw on theories of content but consider whether LLMs can be ascribed propositional attitudes, and Stoljar & Zhang (2025) for a discussion of whether LLMs can think.
+[^4] See goldsteinDoesChatGPTHave2024 and chalmersPropositionalInterpretabilityArtificial2025 for discussions that also draw on theories of content but consider whether LLMs can be ascribed propositional attitudes, and stoljarWhyChatGPTDoesnt2025 for a discussion of whether LLMs can think.
 We do not claim that these models ‘know’ what they mean, or even what meaning is, nor that they have any conception of truth or falsity.
 Our sole contention is that the problem of grounding internal states and outputs in the world is solvable in principle, and in some cases arguably already in practice, for LLMs and related models.
-While having grounded internal states may be a crucial step toward understanding, agency, knowledge, and various other cognitive and mental capacities, it is unlikely that they are, by themselves, sufficient for those capacities.
-This clarification is especially important, as the current debate occasionally blurs these distinctions – a conflation we strive to avoid in this paper.
-The paper proceeds as follows.
+While having grounded internal states may be a crucial step toward understanding, agency, knowledge, and various other cognitive and mental capacities, it is unlikely that such states are, by themselves, sufficient for those capacities.
+This clarification is especially important, as the current debate occasionally blurs these distinctions—a conflation we strive to avoid in this paper.
+We proceed as follows.
 First, we introduce the Symbol Grounding Problem in the context of classical AI (section 2).
 Next, we present the Vector Grounding Problem as it applies to LLMs and related models (section 3).
-We then differentiate between various senses of grounding, showing that only one of them – referential grounding – is essential to the Vector Grounding Problem (section 4).
-Drawing on philosophical theories of representational content, we articulate a set of conditions that must be fulfilled by a system to achieve referential grounding (section 5).
-Finally, we argue that at least some LLMs and related multimodal models satisfy these requirements, and therefore have internal representations and outputs with intrinsic meaning (section 6).
+We then differentiate between various notions of grounding, showing that only one of them—referential grounding—is essential to the Vector Grounding Problem (section 4).
+Drawing on philosophical theories of representational content, we articulate a set of conditions that a system must satisfy to achieve referential grounding (section 5).
+We argue that at least some LLMs and related multimodal models satisfy these requirements, and therefore have internal representations and outputs with intrinsic meaning (section 6).
+Finally, we discuss some surprising implications of our argument (section 7).
 2 The Classical Symbol Grounding Problem
-Research on artificial intelligence has historically followed two competing paradigms: classical (or symbolic), and connectionist.
-The classical paradigm focuses on systems that represent and manipulate information using discrete symbols, based on explicit rules or procedures.
-Symbolic manipulation is purely syntactic; it is based on the shapes of symbols and the way in which they can be combined into complex symbolic structures (Simon & Newell 1971).
-In contrast, the connectionist paradigm focuses on artificial neural networks that process information in a distributed way through interconnected nodes, learning patterns from data rather than following explicit symbolic rules (Mcclelland et al.
-1986).
-The Symbol Grounding Problem, as identified by Harnad (1990), asks how the meaning of symbols in a formal symbol system can be intrinsic to the system itself, rather than merely parasitic on the meanings in the minds of external interpreters.
+Research in artificial intelligence has historically been shaped by two distinct paradigms: the classical (or symbolic) approach and the connectionist approach.
+The classical paradigm focuses on systems that represent and manipulate information through discrete symbols according to explicit rules.
+This symbolic manipulation is purely syntactic; it depends solely on the formal properties of symbols and the rules that govern their combination into complex structures (simonHumanProblemSolving1971).
+By contrast, the connectionist paradigm uses artificial neural networks that process information through massively parallel and distributed computation across interconnected nodes.
+Rather than following predetermined symbolic rules, these systems induce patterns directly from data through learning algorithms (mcclellandParallelDistributedProcessing1986).
+The Symbol Grounding Problem, as identified by Harnad1990, asks how the meaning of symbols in a formal symbol system can be intrinsic to the system itself, rather than merely parasitic on the meanings in the minds of external interpreters.
 More precisely:
 Symbol Grounding Problem.
-How can symbol tokens, manipulated solely on the basis of their shapes according to syntactic rules, acquire meaning and refer to entities and properties in the real world – rather than merely relating other meaningless symbol tokens – without depending on the semantic interpretations provided by external meaning-bearing systems?
-[^4] It is common to take Harnad’s Symbol Grounding Problem as an elaboration of an earlier, highly influential criticism of AI: Searle (1980)’s Chinese Room thought experiment.
-Harnad (1990) himself followed this line.
-Harnad (1990) provides a simple and powerful intuition pump to motivate the existence of a Symbol Grounding Problem in classical, symbolic AI: the Chinese Dictionary.
-Take a system, be it biological or artificial, which knows no natural language.
-Suppose that we want it to learn Chinese (or any other language), but the only resource made available to the system is a single-language dictionary, say a comprehensive Chinese-Chinese dictionary.
-Will the system be able to learn Chinese just by using the dictionary?
-The thought experiment invites the intuition that the system would not be able to do so.
-After all, the most it can do is to look up strings of symbols and see what other strings of symbols follow them.
-The former would be the words to be defined, the latter their definitions.
-However, by themselves the symbols are just meaningless shapes followed by other shapes, and it seems intuitively hopeless that the meaning of words (parts of those strings of symbols) could be learned this way.
-What seems to be missing is a way out from the symbolic ‘merry-go-round’: to connect those symbols in the dictionary to what they pick out in the world.
-But how can that be possible if the system has no access to the world, but only to further symbols?
-Symbolic AI systems, according to Harnad, find themselves in this predicament.
-All they have access to are strings of symbols and how they follow each other.
-There seems to be no way for connections between such symbols and the world to be established, and thus no way for the symbols to have intrinsic meaning.
-Humans, who do have representations arguably acquired by means of engagement with the world, are, on the other hand, able to interpret those symbolic representations and outputs as meaningful.
-The meaning of the representations is hence extrinsic, depending on human interpreters.
-Take out the interpreter, and no meaning is to be found.
-In brief, symbolic AI systems, the argument goes, manipulate symbols that are not connected to the world, that are not grounded in the world; and thereby have no intrinsic meaning.
-Classical AI systems, in other words, suffer from the Symbol Grounding Problem.
+How can symbol tokens, manipulated solely on the basis of their shapes according to syntactic rules, acquire meaning and refer to entities and properties in the real world—rather than merely relating to other meaningless symbol tokens—without depending on the semantic interpretations provided by external meaning-bearing systems?
+[^5] It is common to take Harnad’s Symbol Grounding Problem as an elaboration of an earlier, highly influential criticism of AI: Searle1980’s Chinese Room thought experiment.
+Harnad1990 himself followed this line.
+Harnad1990 provides a compelling intuition pump to motivate the Symbol Grounding Problem in classical symbolic AI.
+Consider a system—biological or artificial—with no prior knowledge of any natural language.
+Suppose we attempt to teach this system Chinese (or any other language), but provide it with only a monolingual resource: a comprehensive Chinese-Chinese dictionary.
+Can the system learn Chinese from this dictionary alone?
+The thought experiment invites a negative answer.
+The system can look up strings of symbols and find other strings that purportedly define them—words paired with their definitions.
+However, these symbols remain mere meaningless shapes followed by other shapes.
+Without any connection to the external world, it seems impossible for the system to grasp the actual meaning of these words.
+The system is trapped in a symbolic ‘merry-go-round’, endlessly cycling through symbols that refer only to other symbols, never breaking free to establish connections with real-world referents.
+Without access to anything beyond symbols, how can the system ever ground these symbols in the real world?
+According to Harnad, symbolic AI systems find themselves in this predicament.
+Limited to manipulating strings of symbols and their syntactic relationships, these systems lack any mechanism for establishing connections between symbols and their real-world referents.
+Consequently, the symbols possess no intrinsic meaning.
+While humans—who do have representations acquired through direct engagement with the world—can interpret these symbolic outputs as meaningful, this meaning remains extrinsic, entirely dependent on human interpretation.
+Remove the human interpreter, and the symbols become meaningless shapes.
+In essence, the argument shows that symbolic AI systems manipulate ungrounded symbols: symbols disconnected from the world they purport to represent.
+These systems thus lack intrinsic semantic content.
+Classical AI systems, in other words, are fundamentally limited by the Symbol Grounding Problem.
 3 The Vector Grounding Problem
-Connectionist models have come a long way since the 1980s, with deep neural networks (DNNs) emerging as powerful tools for various tasks (LeCun et al.
-2015).
-Recent research has converged onto training deep neural networks on vast amounts of data to perform a broad range of tasks they were not explicitly designed for.
-Such DNNs are routinely called ‘Foundation Models’ (Bommasani et al.
-2022), though we will use the more neutral label ‘Generative Pre-Trained Models’.
-Many among the most popular Generative Pre-Trained Models today are powered by the Transformer architecture, first introduced by Vaswani et al.
-(2017).
-Transformers use an attention mechanism to model global dependencies between sequential elements in the input, such as sentences in natural language.
-The general inductive biases supplied by this mechanism make Transformers suitable for many domains, including natural language processing and computer vision.
-Some Generative Pre-Trained Models, called Large Language Models (LLMs), are trained exclusively on text.
-State-of-the-art LLMs like GPT-4  and Claude have over $10^{11}$ parameters and are trained on massive corpora spanning hundred of billions of words, containing a significant subset of all written text on the internet.
+Connectionist models have come a long way since the 1980s, with deep neural networks (DNNs) emerging as powerful tools for various tasks (lecunDeepLearning2015).
+Recent research has converged on training deep neural networks on vast amounts of data to perform a broad range of tasks they were not explicitly designed for.
+Such DNNs are often called ‘foundation models’ (bommasaniOpportunitiesRisksFoundation2022), though we will use the more descriptive label ‘generative pre-trained models’.
+Most of these models are built with a neural network architecture called the ‘Transformer’, introduced by Vaswani2017, which uses a mechanism called ‘attention’ to capture global dependencies between sequential elements.
+This architecture has proved particularly effective for linguistic and visual data.
+LLMs are generative pre-trained models trained on text data.
+State-of-the-art LLMs have over $10^{11}$ parameters and are trained on massive corpora spanning trillions of words, effectively ingesting a substantial fraction of all text available on the internet.
 After training, LLMs can generate well-written and coherent paragraphs about virtually any topic.
-Furthermore, given explicit instructions in natural language in the input sequence, or ‘prompts’, they can perform a broad range of tasks, including: summarising long-form content, translating between multiple languages, answering complex questions, engaging in commonsense reasoning, solving simple logic and maths problems, generating code for simple programs, or explaining jokes (Srivastava et al.
-2023).
-LLMs process input text into a sequence of tokens (i.e., full words, word-parts such as ‘-ing’, punctuation marks, and other linguistic symbols).
-LLMs encode each token as a vector of real-valued numbers – known as an embedding – in a high-dimensional vector space, and estimate the probabilities of all possible next tokens in the sequence based on the previous embeddings.
-The resulting prediction of the next token in the sequence is then compared to the actual next token, and the model’s parameters are subsequently updated to make the correct token more probable in that context, and other tokens less probable.
-While LLMs do not treat words as the main units of processing, the combination of tokens can form an indefinite number of possible words, and we can view the vectors for such composing tokens as jointly forming a word embedding.
+Furthermore, given explicit instructions in natural language in the input sequence, or ‘prompts’, they can perform a broad range of tasks, including: summarising long-form content, translating between multiple languages, answering complex questions, engaging in commonsense reasoning, solving logic and maths problems, generating code, or explaining jokes (srivastava2023beyond).
+LLMs process text into a sequence of tokens: units that may map onto whole words or word parts.
+Each token is encoded as a high-dimensional vector called an embedding.
+The model then estimates a probability distribution over possible next tokens based on the sequence of token embeddings.
+During training, next-token predictions are compared against the actual next token in the training data, and model parameters are adjusted to increase the probability of correct tokens while suppressing alternatives.
+Though LLMs do not treat words as the main units of processing, the combination of tokens can form an indefinite number of possible words, and we can view the vectors for such composing tokens as jointly forming a word embedding.
 This training process (confusingly known as pre-training) encourages the model to produce context-sensitive embeddings, with semantically similar words having vectors that are close together in the high-dimensional space.
-There is a deep connection between the way in which LLMs learn embeddings from linguistic corpora, and the old idea from structural linguistics according to which word meaning can be inferred from the contexts in which they appear, known as the ‘distributional hypothesis’ (Harris 1954, Firth 1957).
-In LLMs, semantic similarity is determined based on distribution across contexts: words that appear in similar contexts are modelled by nearby vectors.
-The dimensions of the vectors are tuned by the model to best predict surrounding words, and capture distributional patterns.
-Although individual dimensions are not typically interpretable, the positions of vectors in the space are a result of statistical patterns in the linguistic data, and reflect semantic similarities in light of how the words are used in context.
-[^5] Vision-Language Models (VLM) that can process images in addition to text work in a similar way: they encode text and images as vectors in a single high-dimensional space.
-We will discuss whether these models have specific implications for the grounding problem in subsection 7.3.
-In this way, embeddings in LLMs are not arbitrary or conventional – as the symbols in classical AI systems typically are – since they are shaped by statistical patterns in language.
-However, LLMs appear to face a similar predicament to that of traditional symbolic AI systems when it comes to grounding and intrinsic meaning.
-Their reliance on continuous vectors, rather than discrete symbols, seems to make no difference to escaping the merry-go-round of representations, and being connected to the external world – LLMs are trapped in a carousel of numbers, rather than symbols.
-Much like their classical counterparts, LLMs manipulate vectors that do not seem grounded in the world and thus lack intrinsic meaning.
-Consequently, they grapple with an analogous difficulty, which we call the Vector Grounding Problem.
-Mirroring Harnad’s 1990 Symbol Grounding Problem, the Vector Grounding Problem asks how the vector embeddings inside Generative Pre-Trained Models can have intrinsic meanings, rather than merely being parasitic on the meanings in the minds of external interpreters.
-More precisely:
+There is a deep connection between the way in which LLMs learn embeddings from linguistic corpora, and the old idea from structural linguistics according to which word meaning can be inferred from the contexts in which they appear, known as the distributional hypothesis (harrisDistributionalStructure1954, firthPapersLinguistics193419511957).
+In LLMs, semantic similarity is determined based on distribution across contexts: words that appear in similar contexts are mapped to nearby vectors.
+Although individual dimensions of the vector space are not typically interpretable, the geometric relationships between vectors encode statistical regularities that reflect latent semantic structure.
+Consequently, embeddings are not purely conventional, unlike symbols in classical AI systems, as their relational structure is determined by statistical patterns in linguistic data.
+[^6] Vision-Language Models (VLMs) extend this approach by encoding both textual and visual inputs within a unified vector space.
+We examine the implications of such multimodal architectures for the grounding problem in subsection 7.3.
+Nevertheless, LLMs seem to face a challenge analogous to that of traditional symbolic AI systems regarding grounding and intrinsic meaning.
+Their use of continuous vectors rather than discrete symbols does not help them escape the ‘merry-go-round’ of representations or connect to the external world—LLMs seem trapped in a carousel of numbers instead of symbols.
+Like their classical counterparts, LLMs manipulate vectors that appear to lack grounding in the world and thus have no intrinsic meaning.
+Consequently, they encounter an analogous predicament, which we call the Vector Grounding Problem:
 Vector Grounding Problem.
-How can vector embeddings, manipulated solely based on their numerical values according to algebraic operations, acquire meaning and refer to entities and properties in the real world – rather than merely relating other meaningless vectors – without depending on the semantic interpretations provided by external meaning-bearing systems?
-Both Grounding Problems, as stated here, focus on internal states, such as symbols and vector embeddings.
-If the Grounding Problems are solved, and such internal states come to have intrinsic meaning, we get straightforward reasons to think that the systems in question produce outputs that are intrinsically meaningful.
-For the meaning of the outputs would thus derive from the meaning of the system’s internal representations, rather than from the interpretations given by human users.
-[^6] In such a case, human users could misinterpret the system’s output if their interpretation does not conform to what the system actually represented – in a way akin to how humans can misinterpret each other.
-It is worth noting that the Grounding Problems are particularly relevant to systems that generate novel outputs.
+How can vector embeddings, manipulated solely on the basis of their numerical values according to algebraic operations, acquire meaning and refer to entities and properties in the real world—rather than merely relating to other meaningless vectors—without depending on the semantic interpretations provided by external meaning-bearing systems?
+Both versions of the Grounding Problem, as stated here, focus on the semantic content of internal states such as symbols and vector embeddings.
+If either problem is solved, so that their internal states acquire intrinsic meaning, we would have compelling grounds for attributing genuine semantic content to the system’s outputs: the meaning of outputs would derive from the system’s internal representations, rather than from post-hoc interpretations by human users.
+[^7] This would create the possibility of genuine misinterpretation: human users could misconstrue the system’s output by failing to grasp what the system actually represented, analogous to how humans routinely misinterpret each other’s statements.
+The Grounding Problems are particularly relevant to systems that generate novel outputs.
 This sets them apart from systems that merely reproduce outputs generated by humans, such as paper copiers.
 When a human-generated text is copied, there is no question about its intrinsic meaning: it was produced by a system with meaningful internal representations to convey a specific meaning to the reader.
-In contrast, generative systems like LLMs produce outputs that are not, for the most part, mere copies of human-generated text; they generate entirely novel sentences whose meanings cannot be equated to what any specific human has ever thought or said.
+In contrast, generative systems like LLMs produce outputs that are not, for the most part, mere copies of human-generated text; they generate novel sentences whose meanings cannot be equated to what any specific human has ever thought or said.
 Unlike with paper copiers, the question of whether LLM outputs are meaningful cannot be straightforwardly explained by appealing to a human author.
-[^7] We thank an anonymous referee for pressing us on this point.
-Bender & Koller (2020) propose a thought-provoking thought experiment, the “Octopus Test”, to pump intuitions about the inescapability of the Vector Grounding Problem for LLMs:
-The Octopus Test.
+[^8] We will return to the question of whether the representations of LLMs only acquire extra-linguistic content through artefactual functions that depend on the intentions of their human designers in section 6.2.1.
+Bender2020 propose a thought experiment, the ‘Octopus Test’, to pump intuitions about the inescapability of the Vector Grounding Problem for LLMs:
+Octopus Test.
 Two English-speaking castaways find themselves stranded on neighbouring islands, separated by treacherous waters.
-Fortuitously, they discover telegraphs left by previous inhabitants, connected via an underwater cable, which enables them to communicate by exchanging telegraphic messages.
-Unbeknownst to them, a superintelligent octopus inhabits these waters and taps into the underwater cable, intercepting their messages.
-Though the octopus lacks any knowledge of English, its superintelligence allows it to detect statistical patterns in the telegraphic messages and to form an accurate representation of the statistical relationships between various telegraphic signals.
-The octopus decides to sever the underwater cable and position itself at the ends of the two resulting cable segments, receiving and replying to telegraphic signals from both castaways based on the statistical patterns it has identified.
-Whether or not the castaways notice this change, the messages sent by the octopus intuitively seem to hold no intrinsic meaning.
-After all, the octopus simply adheres to the statistical patterns it has learned from listening in on the previous exchanges between the humans, without any understanding of the human interpretation of the signals, such as ‘coconut’ or ‘tree’.
-Furthermore, the octopus likely does not comprehend that the signals possess meaning or serve a communicative function.
-[^8] Bender & Koller (2020) emphasise the fact that the octopus lacks communicative intent when generating telegraphic signals.
-Since our focus here is the narrower, more basic question of vector grounding, rather than the more complex capacities related to language comprehension and communication, we will set those considerations aside.
-As should be clear, the octopus in this thought experiment is operating roughly as an extremely powerful LLM would.
-Regardless of how the octopus represents the signals and their statistical relations, be it through vectors or other means, it only has access to the signals and their statistical relations – without any connection to the entities and properties in the real world that the words encoded in those signals represent.
-Consequently, even if the octopus can produce outputs that appear sensible to humans in all situations, these outputs lack intrinsic meaning.
-[^9] Our version of the Octopus Test goes beyond Bender & Koller (2020) by assuming that it is possible the humans would never notice they were no longer communicating with each other.
-In contrast, Bender & Koller (2020) contend that the octopus would fail to produce appropriate outputs to questions involving concrete objects it has never perceived, such as ‘how do I build a coconut catapult?’.
-This claim seems to be challenged by the often plausible replies that current LLMs provide to this sort of question.
-We should expect the superintelligent octopus to generate even more reliably plausible responses.
-The Octopus Test implies that LLMs generate outputs as devoid of intrinsic meaning as those produced by the superintelligent octopus.
-[^10] Arguably, LLMs’ outputs would be even more devoid of intrinsic meaning than the octopus’, given that the latter is a biological organism with cognitive states, motivation, superintelligence, and other features that LLMs lack.
-Similar considerations apply to the comparison between LLMs and parrots (Bender et al.
-2021).
+Fortuitously, they discover telegraphs left by previous inhabitants, connected via an underwater cable, enabling them to communicate through telegraphic messages.
+Unbeknownst to them, a superintelligent octopus inhabiting these waters taps into the cable, intercepting their messages.
+Though the octopus has no knowledge of English, its superintelligence enables it to detect statistical patterns in the telegraphic exchanges and construct an accurate model of the statistical relationships between various signals.
+The octopus then severs the cable and positions itself at both ends, intercepting messages from each castaway and generating responses based solely on the statistical patterns it has learned.
+Whether or not the castaways detect this substitution, the octopus’s messages intuitively seem devoid of intrinsic meaning.
+After all, the octopus merely reproduces statistical patterns gleaned from eavesdropping on human exchanges, without understanding what the signals denote.
+Moreover, the octopus likely fails to grasp that these signals carry meaning or serve any communicative function.
+[^9] Bender2020 emphasize that the octopus lacks communicative intent when generating telegraphic signals.
+Since our focus is on the narrower question of vector grounding rather than the broader capacities underlying language comprehension and communication, we set aside these considerations.
+The octopus in this thought experiment operates analogously to a powerful LLM.
+Regardless of the format of its internal states—whether vectors or otherwise—the octopus accesses only linguistic signals and their statistical relationships, lacking any connection to the real-world entities and properties these signals purportedly represent.
+Consequently, even if the octopus produces outputs that appear perfectly sensible to human observers, these outputs lack intrinsic meaning.
+The Octopus Test suggests that LLMs generate outputs as devoid of intrinsic meaning as those produced by the superintelligent octopus.
+[^10] One might argue that LLMs’ outputs are even more devoid of intrinsic meaning than the octopus’s, given that the latter is a biological organism endowed with cognitive states, motivations, superintelligence, and other features that LLMs lack.
+Similar considerations apply when comparing LLMs to parrots (Bender2021).
 In the case of these animals, even if their outputs carry some meaning, they typically do not align with the meaning of the language they produce.
-It is humans, whether castaways or otherwise, who interpret these outputs as meaningful because they resemble meaningful words and sentences.
-LLMs have no access to coconuts, trees, or anything else beyond language.
-The vectors they receive and manipulate during internal processing are merely meaningless arrays of numbers, ungrounded in anything outside statistical patterns in language, just like the outputs they generate.
-In other words, the Octopus Test invites the intuition that all meaning in the outputs produced by LLMs is extrinsic, i.e., it lies in the eye of the beholder, their human interpreters.
-This would be so, we might add, because LLMs are susceptible to the Vector Grounding Problem.
-Throughout the remainder of this paper, we will challenge this claim and counter the intuition primed by the Octopus Test.
-We aim to show that, contrary to first appearances, LLMs indeed possess the features required to generate meaningful outputs based on meaningful internal representations.
-Before developing this argument further, it is essential to clarify the target notion of grounding.
+Humans (castaways or contemporary users) project meaning onto these outputs because they superficially resemble meaningful words and sentences.
+LLMs have no access to coconuts, trees, or any extra-linguistic reality.
+The vectors they process are merely arrays of numbers, grounded only in statistical patterns within language, not in the world these patterns purportedly describe.
+In other words, the Octopus Test invites the conclusion that all meaning in the outputs produced by LLMs is extrinsic: it lies in the eye of the beholder.
+This would be because LLMs are susceptible to the Vector Grounding Problem.
+Throughout the remainder of this paper, we challenge this claim and the intuitions primed by the Octopus Test.
+We argue that, contrary to initial appearances, LLMs possess the necessary features to generate meaningful outputs based on internal representations with extra-linguistic content.
+Before developing this argument, we must first clarify our target notion of grounding.
 4 Five notions of Grounding
@@
 With this in mind, let us examine each notion of grounding in turn.
+4.1 Referential Grounding
+The first notion of grounding, which we call referential grounding, closely aligns with the notion of reference.
+It is the type of grounding that connects representations to things in the world.
+Reference can be loosely described as a relation that enables representations to ‘hook onto’ worldly entities or properties.
+[^11] Much philosophical work has been dedicated to examining the nature and varieties of reference.
+For our limited purposes, however, the intuitive notion captured by the ‘hooking’ metaphor will suffice.
+For example, the word ‘Paris’ refers to the city of Paris, and arguably, so does a map of Paris.
+Both representations allow us to receive and convey information about the actual city of Paris.
+Similarly, many of our thoughts seem to hook onto specific entities and properties in the world, allowing us to entertain thoughts that refer to Paris, such as beliefs about it being the capital of France or desires to visit the city.
+Referential grounding, as we use the term, is whatever makes it so that representations, whether public (e.g., words and maps) or internal (e.g., beliefs and subpersonal states), hook onto the world.
+It is plausible that different kinds of representation require different means for hooking onto the world in the relevant sense, that is, to be referentially grounded in specific worldly entities and properties (CoelhoMollo2015).
+4.2 Sensorimotor Grounding
+Another type of grounding that is often invoked, particularly in cognitive psychology, is what we might call sensorimotor grounding.
+More precisely, we can divide sensorimotor grounding into two related but different views that are often conflated.
+[^12] We thank an anonymous reviewer for pressing this point.
+According to what we may call the ‘sensorimotor representation’ view of grounding, for an internal representation to be sensorimotorically grounded is for it to be fundamentally linked to sensorimotor representations available within a cognitive system.
+For instance, the lexical representation ‘football’ is sensorimotorically grounded if its representation in a cognitive system can be reduced to, or is intrinsically connected to, sensorimotor representations of kicking movements, of the visual perception of a football pitch, of the auditory perception of the sound of a ball being kicked, etc.
+This kind of grounding plays a prominent role in theories in cognitive psychology that embrace a moderate form of embodied cognition, such as those positing the existence of “perceptual symbol systems” in human cognition (Barsalou1999, Pulvermueller1999).
+This kind of sensorimotor grounding is the approach proposed by Harnad1990 to tackle the Symbol Grounding Problem.
+Another possible way to understand sensorimotor grounding is in terms of what we may call the ‘sensorimotor contact’ view of grounding.
+Its guiding idea is that an organism’s sensory and motor systems interact with the world in a distinctive, unmediated way.
+Specifically, sensory and motor transducers—such as nerve endings in sensory surfaces—make direct contact with the world, thus providing a basic form of grounding that is inherited, through appropriate causal chains, by sensorimotor systems and organisms as a whole.
+4.3 Relational Grounding
+A third notion of grounding, that we may call relational grounding, is occasionally used to refer specifically to intra-linguistic relationships, in accordance with the idea that a word’s meaning in a language is partly determined by its relations to other words.
+[^13] The label is somewhat under-descriptive, since all forms of grounding are relational, though in different ways.
+What we wish to underline, however, is that on this sense of grounding the meaning of representations depends on the relations they stand in to other representations in the same representational system.
+This “language-to-language grounding” (parteeMontagueGrammarMental1981) captures the way in which a language learner can acquire knowledge about the meanings of words by learning semantic connections between words in the language.
+Relational grounding is closely related to inferential role semantics (blockAdvertisementSemanticsPsychology1986, brandomMakingItExplicit1994), which identifies the meaning of a linguistic expression with its inferential relations to other expressions.
+There is also a clear connection between relational grounding and distributional semantics.
+Distributional semantics seeks to analyse word meanings in terms of word-to-word relationships.
+However, instead of drawing on linguistic analyses of word meaning, it relies on analysing co-occurrence statistics in text corpora.
+Consequently, distributional semantic models capture relationships between words grounded in patterns of use, rather than linguistic theory.
+The high dimensionality of vector spaces in LLMs makes it possible to represent complex, multifaceted similarities between words—some of which plausibly reflect conceptual structure and/or inferential roles (Piantadosi2022).
+4.4 Communicative Grounding
+The notion of grounding is also used in relation to another aspect of language, having to do with communication (clarkGroundingCommunication1991, traumComputationalTheoryGrounding1994).
+In this context, the ‘grounding problem’ refers to the problem of establishing common ground between speakers in conversation.
+From this perspective, linguistic communication can be seen as a form of coordinated action that involves collaborating to reach a common understanding of what is said.
+The process of communicative grounding may involve explicit clarification strategies, such as prompting an interlocutor to repeat or reformulate a garbled or ambiguous statement.
+It also involves tracking communicative intentions during the exchange, to infer what is communicated from what is said.
+Such inferences can be made partly on the basis of pragmatic presuppositions, namely, what speakers take for granted or assume to be true when they use certain sentences in the context of the conversation (stalnakerCommonGround2002).
+The communicative notion of grounding has also been discussed by natural language processing (NLP) researchers (brennan1998grounding, dimaroComputationalGroundingOverview2021, chanduGroundingGroundingNLP2021), and is closely tied to the challenge to LLM meaning advanced by Bender2020, who define meaning as a relation between linguistic expressions and communicative intent, characterising this relation as a form of grounding.
+While communicative and referential grounding are distinct notions, they are related in important ways.
+Referential grounding anchors language to the world, but the same linguistic expressions may be referentially grounded in different ways for different speakers, due to differences in perception, knowledge, and conceptualisation.
+This means that calibrating reference in conversation requires communicative grounding—language users must work together to negotiate a common way of connecting language to the world.
+Without such collaboration, there is no guarantee that speakers really mean the same thing, even if they are using the same words.
+4.5 Epistemic Grounding
+Finally, the notion of grounding is occasionally used by NLP researchers to refer more minimally to the relationship between linguistic expressions and information stored in knowledge bases (poonGroundedUnsupervisedSemantic2013, tsaiConceptGroundingMultiple2016, chanduGroundingGroundingNLP2021, thoppilanLaMDALanguageModels2022).
+For example, an expert system may contain a knowledge base of facts and rules belonging to a particular domain.
+When the system receives user input, it can query this knowledge base to determine an appropriate response.
+A classic example is MYCIN, an expert system for diagnosing infectious blood diseases (shortliffeMycinKnowledgeBasedComputer1977).
+When given patient symptoms, MYCIN would search its knowledge base of medical rules and facts to suggest possible diagnoses and recommend treatments.
+In this system, language is grounded in the structured knowledge contained in the expert system’s knowledge base.
+We call this notion epistemic grounding.
+There are two ways in which current LLMs may involve epistemic grounding.
+Firstly, they can be hooked up to external databases used to retrieve information (borgeaudImprovingLanguageModels2022, thoppilanLaMDALanguageModels2022).
+Secondly, it has been suggested that even LLMs that are not augmented with retrieval capabilities might be functionally similar to knowledge bases themselves (petroniLanguageModelsKnowledge2019).
+Indeed, they may encode a significant amount of world knowledge in their weights after pre-training, as illustrated by the remarkable performance of state-of-the-art LLMs on a wide range of factual question-answering benchmarks (openaiGPT4TechnicalReport2023, bubeckSparksArtificialGeneral2023).
 Figure 1: Five notions of grounding.
@@
 Epistemic grounding: A lexical representation (<DOG>) is connected to information stored in a knowledge base (about dog facts).
-4.1 Referential Grounding
-The first notion of grounding, which we call referential grounding, closely aligns with the philosophical notion of reference.
-[^11] Partee (1981) uses the phrase “language-to-world grounding” to designate a closely related notion, which she contrasts with “language-to-language grounding” (see subsection 4.3 below for discussion of the latter).
-While Partee’s notion of language-to-world grounding focuses on what ties down the intensions of linguistic expressions with respect to their extra-linguistic content, our notion of referential grounding is broader, including the relation between internal representations and their worldly referents.
-It is the type of grounding that connects representations to things in the world.
-Reference can be loosely described as a relation that enables representations to ‘hook onto’ worldly entities or properties.
-[^12] Much philosophical work has been dedicated to examining the nature and varieties of reference.
-For our limited purposes, however, the intuitive notion captured by the ‘hooking’ metaphor will suffice.
-For example, the word ‘Paris’ refers to the city of Paris, and arguably, so does a map of Paris.
-Both representations allow us to receive and convey information about the actual city of Paris.
-Similarly, many of our thoughts seem to hook onto specific entities and properties in the world, allowing us to entertain thoughts that refer to Paris, such as beliefs about it being the capital of France or desires to visit the city.
-Referential grounding, as we use the term, is whatever makes it so that representations, whether public (e.g., words and maps) or internal (e.g., beliefs and subpersonal states), hook onto the world.
-It is plausible that different kinds of representation require different means for hooking onto the world in the relevant sense, that is, to be referentially grounded in specific worldly entities and properties (Coelho Mollo 2015).
-4.2 Sensorimotor Grounding
-Another type of grounding that is often appealed to, especially in cognitive psychology, is what we call sensorimotor grounding.
-For an internal representation to be grounded in this sense is for it to be fundamentally linked to sensorimotor representations available within a cognitive system.
-For instance, the lexical representation ‘football’ is sensorimotorically grounded if its representation in a cognitive system can be reduced to, or is intrinsically connected to, sensorimotor representations of kicking movements, of the visual perception of a football pitch, of the auditory perception of the sound of a ball being kicked, etc.
-This kind of grounding plays a prominent role in theories in cognitive psychology that embrace a moderate form of embodied cognition, such as those positing the existence of “perceptual symbol systems” in human cognition (Barsalou 1999, Pulvermüller 1999).
-Interestingly, sensorimotor grounding is also the solution proposed by Harnad (1990) to his Symbol Grounding Problem.
-He argued that the symbols in classical AI systems needed to be grounded in two types of representation: (a) iconic representations, which depict transformations of the projections of object ‘shapes’ on a system’s sensory surfaces; and (b) categorical representations, which represent invariant sensory features that facilitate the identification and discrimination of objects in different categories.
-Harnad hypothesised that connectionist approaches to AI would be better equipped to provide the necessary iconic and categorical representations, which could then complement symbolic AI systems and result in hybrid symbolic-connectionist systems.
-We will come back to Harnad’s suggestion in subsection 4.6, arguing that it fails to solve the Symbol Grounding Problem, and does not help with the Vector Grounding Problem either.
-4.3 Relational Grounding
-A third notion of grounding, that we may call relational grounding, is occasionally used to refer specifically to intra-linguistic relationships, in accordance with the idea that a word’s meaning in a language is partly determined by its relations to other words.
-[^13] The label is somewhat under-descriptive, since all forms of grounding are relational, though in different ways.
-What we wish to underline, however, is that on this sense of grounding the meaning of representations depends on the relations they stand in to other representations in the same representational system.
-Partee (1981) introduces the notion of “language-to-language grounding”, referring to the way in which a language learner acquires knowledge about the meanings of words by learning semantic connections between words in the language.
-According to Partee, one key way in which this happens is through accepting certain sentences as true, particularly those that are regarded as impervious to empirical challenge within the linguistic community.
-For example, learning that “bachelors are unmarried males” helps ground the meaning of ‘bachelor’ through its relation to ‘unmarried’ and ‘male’.
-Relational grounding is closely related to inferential role semantics (Block 1986, Brandom 1994).
-Inferential role semantics identifies the meaning of a linguistic expression with its inferential relations to other expressions.
-For example, the word ‘and’ plays a logical role in reasoning that determines when compound statements are true or false based on the truth or falsity of their components.
-Likewise, words like ‘larger’ or ‘slower’ have inferential roles in comparisons, allowing us to determine comparisons between items or the relative properties of a single item.
-Knowledge of such relations between words underlies our ability to draw material inferences in language and thought.
-Finally, there is a clear connection between Partee’s “language-to-language grounding” and distributional semantics.
-Distributional semantics seeks to analyse word meanings in terms of word-to-word relationships.
-However, instead of drawing on linguistic analyses of word meaning, it relies on analysing co-occurrence statistics in text corpora.
-Consequently, distributional semantic models capture relationships between words grounded in patterns of use, rather than linguistic theory.
-The high dimensionality of vector spaces in LLMs makes it possible to represent complex, multifaceted similarities between words – some of which plausibly reflect conceptual structure and/or inferential roles (Piantadosi & Hill 2022).
-4.4 Communicative Grounding
-The notion of grounding is also used in relation to another aspect of language, which has to do with communication (Clark & Brennan 1991, Traum 1994).
-In this context, the ‘grounding problem’ refers to the problem of establishing common ground between speakers in conversation.
-From this perspective, linguistic communication can be seen as a form of coordinated action that involves collaborating to reach a common understanding of what is said.
-The process of communicative grounding may involve explicit clarification strategies, such as prompting an interlocutor to repeat or reformulate a garbled or ambiguous statement.
-It also involves tracking communicative intentions during the exchange, to infer what is communicated from what is said.
-Such inferences can be made partly on the basis of pragmatic presuppositions, namely, what speakers take for granted or assume to be true when they use certain sentences in the context of the conversation (Stalnaker 2002).
-The communicative notion of grounding has also been discussed by natural language processing (NLP) researchers (Brennan 1998, Di Maro 2021, Chandu et al.
-2021).
-For example, Chandu et al.
-(2021) distinguish between static grounding, in which common ground in basic communication between a human user and a virtual agent is automatically established by querying a database (e.g., asking for a weather report); and dynamic grounding, in which common ground is more gradually established in back-and-forth communication between a human and a dialogue agent.
-Bender & Koller (2020) define meaning to be a relation between linguistic expressions and communicative intent, and characterise this relation as a form of grounding.
-While communicative and referential grounding are distinct notions, they are related in important ways.
-Referential grounding anchors language to the world, but the same linguistic expressions may be referentially grounded in different ways for different speakers, due to differences in perception, knowledge, and conceptualisation.
-This means that calibrating reference in conversation requires communicative grounding – language users must work together to negotiate a common way of connecting language to the world.
-Without such collaboration, there is no guarantee that speakers really mean the same thing, even if they are using the same words.
-4.5 Epistemic Grounding
-Finally, the notion of grounding is occasionally used by NLP researchers to refer more minimally to the relationship between linguistic expressions and information stored in knowledge bases (Poon 2013, Tsai & Roth 2016, Chandu et al.
-2021, Thoppilan et al.
-2022).
-For example, an expert system may contain a knowledge base of facts and rules belonging to a particular domain.
-When the system receives user input, it can query this knowledge base to determine an appropriate response.
-A classic example is MYCIN, an expert system for diagnosing infectious blood diseases (Shortliffe 1977).
-When given patient symptoms, MYCIN would search its knowledge base of medical rules and facts to suggest possible diagnoses and recommend treatments.
-In this system, language is grounded in the structured knowledge contained in the expert system’s knowledge base.
-We call this notion epistemic grounding.
-There are two ways in which current LLMs may involve epistemic grounding.
-Firstly, they can be hooked up to external databases used to retrieve information (Borgeaud et al.
-2022, Thoppilan et al.
-2022).
-Secondly, it has been suggested that even LLMs that are not augmented with retrieval capabilities might be functionally similar to knowledge bases themselves (Petroni et al.
-2019).
-Indeed, they may encode a significant amount of world knowledge in their weights after pre-training, as illustrated by the remarkable performance of state-of-the-art LLMs on a wide range of factual question-answering benchmarks (OpenAI 2023, Bubeck et al.
-2023).
 4.6 What Grounding for LLMs?
-We have identified five distinct notions of grounding that play prominent roles in discussions about meaning grounding in biological and artificial systems.
-Which of these notions is necessary for LLMs to possess representations and produce outputs endowed with intrinsic meaning?
-(Recall that intrinsic meaning is meaning that is independent from external interpretation.) To identify the most suitable notion of grounding for this task, we will employ an elimination strategy, examining each notion for its adequacy, beginning with Harnad’s1990 proposal of appealing to sensorimotor grounding as a solution to the Symbol Grounding Problem.
-Sensorimotor grounding connects conceptual or linguistic representations to sensorimotor representations within the same cognitive system.
-This sort of grounding can be useful to shed light on some features of the former representations, how they work, and what roles they can play in the economy of a cognitive system.
-However, sensorimotor grounding falls short in addressing the merry-go-round problem, as it merely grounds one type of representation in another: it does not tackle the question of how sensorimotor representations establish a connection with the world.
-One might argue that sensorimotor representations are grounded in the world by resembling, to some extent, the entities they represent.
-However, the success of such an approach relies on an additional factor beyond sensorimotor grounding itself: it is the relevant resemblance relations that are supposed to ground those sensorimotor representations in the world.
-As sensorimotor grounding does not provide a connection with the world but only with other representations, it does not address the Symbol Grounding Problem.
-Similarly, this form of grounding does not resolve the Vector Grounding Problem, as grounding vectors in other vectors still leaves open the question of how any of those vectors come to be connected to the world.
-The same line of reasoning applies to relational, communicative, and epistemic grounding.
-Relational grounding anchors linguistic representations in their relationships to other linguistic representations, suggesting that a representation’s meaning is determined by its connections to the meanings of other representations within a network of inferential, conceptual, or statistical relations.
-Although this kind of grounding clearly does not escape the representational carousel, inferential-role semantics may suffice for connecting at least some words to what they are about, e.g., logical connectives, as their meanings are arguably exhausted by the inferences they are involved in (Block 1986, Fodor 1990).
-Logical connectives are however a rather special case, as they are not supposed to represent entities in the world, but rather abstract relations between any possible entities.
+We have identified five distinct notions that play prominent roles in discussions about grounding in biological and artificial systems.
+Which of these is necessary for LLMs to possess representations and produce outputs with intrinsic meaning?
+(Recall that intrinsic meaning is meaning that is independent from external interpretation.) To identify the most suitable notion, we will proceed by elimination, examining each candidate’s adequacy—beginning with the two kinds of sensorimotor grounding.
+The ‘sensorimotor contact’ version of sensorimotor grounding claims that causal contact between sensorimotor transducers and the world grounds downstream representations.
+While promising, this view does not explain how the brute impinging of light rays, sound waves and the like on sensory surfaces can ground representations in the specific worldly things they are about.
+Indeed, such representations are typically not about light rays or sound waves, but rather about horses, democracies, elms, and so forth.
+[^14] Mutatis mutandis for electric signals that activate muscles vis-à-vis motor plans, actions, and behaviours.
+By itself, this sort of causal contact with the world does not explain how internal representations come to be about the latter, rather than (just) the former.
+As we will see (section 5), an account of how sensorimotor transducers can contribute to grounding requires additional components to be in place, making the sensorimotor contact version of sensorimotor grounding collapse into the notion of referential grounding: the causal contact with the world that sensorimotor transducers allow is just one of the components that can underlie referential grounding.
+As we will argue in section 6, moreover, such unmediated causal contact with the world is not even necessary for referential grounding.
+The other version of sensorimotor grounding, the ‘sensorimotor representation’ view, connects conceptual or linguistic representations to sensorimotor representations within the same cognitive system.
+While this connection can illuminate the relevant representations’ role within a cognitive architecture, it fails to address the merry-go-round problem.
+By grounding one type of representation in another, it merely postpones the crucial question: how do sensorimotor representations themselves establish contact with the world?
+Since sensorimotor representational grounding does not provide a connection with the world but only with other representations, it does not address the Symbol Grounding Problem.
+Similarly, it offers no resolution to the Vector Grounding Problem, as grounding vectors in other vectors leaves unanswered how any vector ultimately connects to worldly entities and properties.
+This line of reasoning extends to relational, communicative, and epistemic grounding.
+Relational grounding anchors linguistic representations to their relationships to other linguistic representations.
+Although this kind of grounding clearly does not escape the representational carousel, it is plausible that it can suffice for endowing at least some representations with intrinsic meaning.
+Consider, for instance, logical connectives, for which an inferential-role semantics is compelling (blockAdvertisementSemanticsPsychology1986, Fodor1990).
+The meaning of a term like ‘and’ is arguably exhausted by the role it plays in licensed inferences; it does not purport to refer to a worldly entity, but to an abstract relation between propositions.
+Insofar as an LLM can induce the correct inferential role for logical terms from its training data, its internal states corresponding to these terms could be said to be meaningfully grounded in that role.
 [^15] Therefore, if an LLM can learn the structure of relations that constitute the meaning of logical connectives and linguistic conjunctions from its training data, then its representations of these terms might be as intrinsically meaningful as they can be.
-Inferential-role semantics can thereby account, at best, for only a very small and special subset of the representations and outputs that biological and artificial systems possess and produce.
-Communicative grounding seeks to anchor representations in the shared content of linguistic interactions between agents.
-However, this shared content relies on the representations within the agents’ cognitive systems and the meanings of any public representations used in the exchange.
-As with sensorimotor grounding, one type of representation is grounded in another, perpetuating the representational carousel without establishing a connection to the world beyond the representations.
-Epistemic grounding, on the other hand, aims to ground linguistic representations in knowledge bases.
-Yet, this approach encounters a familiar obstacle: the representations within the knowledge bases themselves call for an explanation of how they acquire intrinsic meaning.
-Consequently, epistemic grounding alone fails to provide a comprehensive solution to the Grounding Problem, as it does not address the question of how representations come to possess intrinsic meaning to start with.
-In sum, the form of grounding that underlies the other four types and is most relevant to addressing the Grounding Problems is what we have called referential grounding.
-Referential grounding anchors a representation – be it a symbol, a map, or a vector – in the world itself, rather than in other representations.
-It stands as the sole form of grounding that escapes the representational merry-go-round and connects a representation to whatever worldly entity or property it is about.
-[^15] Or almost: see the caveat about inferential-role semantics and logical connectives above.
-Consequently, to determine whether Generative Pre-Trained Models such as LLMs manipulate and produce representations and outputs with intrinsic meaning, we must assess whether they can achieve referential grounding.
-To do so, we first need to examine the requirements for a representation to be referentially grounded.
+However, logical connectives are a rather special case; at best, inferential-role semantics accounts for only a narrow subset of the representations and outputs that biological and artificial systems possess and produce.
+A similar case can be made for the representation of purely syntactic properties.
+The function of an internal state that represents, say, a subject-verb agreement rule or a filler-gap dependency is exhausted by its contribution to generating grammatically well-formed token sequences.
+Its correctness conditions are defined entirely within the formal system of the language itself.
+There is strong empirical evidence that LLMs, through the objective of next-token prediction alone, do in fact develop internal mechanisms that implement such grammatical rules.
+[^16] See hardingOperationalisingRepresentationNatural2023c for a discussion of criteria for ascribing representation of syntactic features to LLMs, and Milliereforthcoming for empirical evidence that LLMs do learn such representations.
+However, the Vector Grounding Problem is not about whether a system can acquire representations of intra-linguistic structures—be they logical or syntactic—but whether it can acquire representations with extra-linguistic content.
+Conceding that LLMs can achieve relational grounding for formal aspects of language does nothing to solve the problem of how their representations could come to be about worldly entities and properties.
+Thus, while relational grounding may well be sufficient for a narrow and special subset of representations whose content is exhausted by their formal role, it remains insufficient for the vast lexicon of terms that purport to refer to the world.
+For these terms, relational grounding merely spins the wheels on the representational merry-go-round; it cannot, by itself, provide the traction needed to connect with extra-linguistic reality.
+In sum, the form of grounding that underlies the others and directly addresses the Grounding Problems is what we have called referential grounding.
+Unlike the alternatives, referential grounding anchors representations in the world, rather than in other representations.
+It alone breaks free from the representational merry-go-round by connecting representations to whatever worldly entity or property they are about.
+Therefore, to determine whether generative pre-trained models like LLMs can manipulate and produce representations with intrinsic meaning, we must assess their capacity for referential grounding.
+This assessment requires examining what conditions must be met for a representation to achieve referential grounding.
 5 Theories of Representational Content: Causes and History
@@
 For instance, ‘edge detectors’ in primary visual cortex (V1) are good candidates for being basic internal representations, whereas ‘shape detectors’ in downstream visual areas are not, since their contents derive partly from the contents of the basic edge detectors.
-Similarly, words and sentences are not basic in this sense, since their meaning derives from other, more basic representational states, such as beliefs, desires and intentions – themselves complex representations.
+Similarly, words and sentences are not basic in this sense, since their meaning derives from other, more basic representational states, such as beliefs, desires and intentions—themselves complex representations.
 The distinction between basic and complex representations cuts across the distinction between intrinsic and extrinsic meaning.
-As a reminder, states with intrinsic meaning are those that carry meaning independently of an interpreter with mental states, such as a human, attributing meaning to them.
 Both the representations in primary visual cortex and the words and sentences humans utter are cases of representation with intrinsic meaning, though the former are basic while the latter are complex.
-By contrast, words randomly formed by the movement of ants in the sand, outputs from a random sentence generator – and, according to Bender & Koller (2020), the outputs of LLMs – only have extrinsic meaning: they are merely shapes that require human interpretation to convey any meaning at all.
-We cannot here do justice to the rich and nuanced interdisciplinary research that has been carried out for the past several decades on the notion of cognitive representation.
-[^16] See Neander (2017), Millikan (2017), Shea (2018) for recent book-length defences of the main contenders today.
+By contrast, words randomly formed by the movement of ants in the sand or outputs from a random sentence generator only have extrinsic meaning: they are merely shapes that require human interpretation to convey any meaning at all.
+We cannot here do justice to the rich and nuanced interdisciplinary research on the notion of cognitive representation.
+[^17] See Neander2017, Millikan2017, Shea2018 for recent book-length defences of the main contenders today.
 What matters for our limited purposes is that, despite some disagreements on the finer details, philosophers of cognitive science have largely reached a consensus on the key factors that enable internal states in a system to become representations, both basic and complex (Shea2018, CoelhoMollo2022).
@@
 Two natural relations are widely considered to be central to endowing basic internal states with meaning: causal-informational relations, and historical relations.
-In general terms, causal-informational relations occur when one state of the world conveys information about another due to a causally-generated correlation between the two states.
+In general terms, causal-informational relations obtain when one state of the world conveys information about another due to a causally-generated correlation between the two states.
 A classic example is the relationship between smoke and fire: smoke is correlated with fire because it is caused by it, so the presence of smoke carries information about the presence of fire.
-Neurons in brain area V1 convey information about edges in the perceived visual scene, as their activation correlates (through a relatively complex causal chain) with the presence of edges in the world in front of the perceiving subject.
-Causal-informational relations capture the main rationale for why systems possess basic internal representations: they allow systems to be sensitive to specific features of their environments – those for which internal states stand in causal-informational relations to.
+Neurons in brain area V1 convey information about edges in the perceived visual scene, as their activation correlates (through a relatively complex causal chain that includes light rays impinging on the retina) with the presence of edges in the world in front of the perceiving subject.
+Causal-informational relations capture the main rationale for why systems possess basic internal representations: they allow systems to be sensitive to specific features of their environments.
 However, the mere existence of a causal-informational relation is insufficient to provide an account of basic representational content.
@@
 In other words, a system must use the presence of smoke to obtain information about the presence of fire, form a belief about there being a fire, guide fire-related behaviour, and so on, for smoke to be a viable candidate for representing fire for that system.
-Similarly, the information about the world that edge detectors in V1 carry into the mammalian brain must be used downstream for further processing – such as detecting food sources – so as to guide appropriate behaviour.
 To be a representation necessarily involves playing a certain role: that of conveying information about the world and being used by the system in virtue of that.
@@
 Causal-informational relations are abundant in nature: every effect carries information about its cause.
-If every such relation should suffice for being a representation, representations would be so extremely trivial that their special explanatory power for cognition would be lost.
-Exploited causal-informational relations alone, however, are still not enough to base an adequate theory of representational content, for another crucial feature of internal representations is missing: their capacity for truth or falsehood, correctness or incorrectness.
-To be a representation inherently involves the possibility of misrepresenting – the possibility of getting things wrong and failing to fit to the world accurately.
+If every such relation should suffice for being a representation, representations would be so trivial that their special explanatory power for cognition would be lost.
+Exploited causal-informational relations alone, however, are still not enough to base an adequate theory of representation, for another crucial feature of internal representations is missing: their capacity for truth or falsehood, correctness or incorrectness.
+[^18] Correctness conditions are typical of so-called descriptive representations.
+Directive representations, e.g., desires and action plans, are in the business of representing how aspects of the world should be.
+Such representations have satisfaction conditions, i.e., the conditions that would satisfy, say, a specific desire.
+For simplicity’s sake, we will be concerned only with descriptive representations.
+To be a representation inherently involves the possibility of misrepresenting.
 V1 neurons may become active when no edge is present, as in the famous Kanizsa’s triangle illusion, thus conveying information to the system that, if not corrected by other subsystems, can lead to inappropriate behaviour (such as trying to touch the non-existent edge).
@@
 They are kinds of things that can be true or false, accurate or inaccurate, appropriate or inappropriate.
-The rationale for systems to possess representations involves that they be accurate enough, enough of the time.
+The rationale for systems to possess representations is that they be accurate enough, enough of the time.
 Otherwise, representations would lead systems astray and hinder appropriate behaviour, rather than the other way around.
-Generally, representations have the function to represent correctly and accurately (to a degree of required precision).
+Generally, (descriptive) representations have the function to represent correctly and accurately (to a degree of required precision).
 [^19] In some special cases, systematic misrepresentation can be a sensible price to pay.
-If the stakes are high enough, it makes sense for representations to err on the side caution: if you are trying to detect predators, it makes sense to produce many false-positive (misrepresenting predators as present when there are none), rather than even one, potentially fatal false-negative.
+If the stakes are high enough, it makes sense for representations to err on the side of caution: if you are trying to detect predators, it makes sense to produce many false-positive (misrepresenting predators as present when there are none), rather than even one, potentially fatal false-negative.
 As the notion of representation is central for cognitive (and computer) science, the notion of function is central to biology.
-This has led philosophers of biology to develop scientifically-grounded, objective theories of function Wright (1973), Millikan (1989), Neander (1991), Godfrey-Smith (1994), Garson (2019).
+This has led philosophers of biology to develop scientifically-grounded, objective theories of function (Wright1973, Millik1989, Neander1991, Godfrey-Smith1994, Garson2019).
 If functions exist out there in the world, they must be a matter of natural features and relations, rather than human interpretation.
-The dominant view of functions is the selected-effects theory Millikan (1989), Neander (1991), Garson (2019).
+The dominant view of functions is the selected-effects theory (Millik1989, Neander1991, Garson2019).
 On this view, functions are ensembles of the historical causes that explain the persistence of a certain trait or feature in a system.
-For instance, hearts today have the function to pump blood because past hearts, by pumping blood, causally contributed to the survival and reproduction of the systems having them.
-Pumping blood is what explains, causal-historically, why hearts still exist – and it is thereby their function.
-The relevant causes can be at the level of the species, typically through natural selection; or at the level of the individual, through selection-like learning processes during development.
-The most influential views on representation incorporate the selected-effects theory to account for internal states having the function to carry information about the world and to guide appropriate behaviour Neander (2017), Millikan (2017), Shea (2018), Coelho Mollo (2022).
-The guiding idea is that, through selection and/or learning, states and processes in cognitive systems come to acquire functions they are supposed to perform.
-For example, neurons in V1 have the function of detecting edges in the visual scene because they were selected for this purpose during evolution via natural selection; and such selection occurred due to the causal-informational relations they maintain with actual edges in the world.
+For instance, hearts today have the function to pump blood because past hearts, by pumping blood, were selected due to their contribution to the survival and reproduction of the systems having them.
+Pumping blood is what explains, through a historical selection process, what hearts were selected for doing—and it is thereby their function.
+The relevant selection processes can be at the level of the species, typically through natural selection; or at the level of the individual, through selection-like learning processes during development.
+The most influential views on representation incorporate the selected-effects theory to account for internal states having the function to carry information about the world and to guide appropriate behaviour (Neander2017, Millikan2017, Shea2018, CoelhoMollo2022).
+The basic idea is that some states and processes in cognitive systems were selected for in virtue of their carrying information about, and guiding successful behaviour toward, states of the world—thus acquiring functions to represent such states.
+For example, neurons in V1 have the function of detecting edges in the visual scene because they were selected by natural selection in virtue of the causal-informational relations they maintain with actual edges in the world.
 When these neurons successfully detect edges, they fulfil their function, representing correctly; when they fail to fire in the presence of an edge or fire in the absence of one, they malfunction, misrepresenting.
-In brief, internal representations are states that carry information about the world, have the function to do so, and are appropriately used by the system in virtue of the information they carry.
-Existing theories of basic representation vary in the emphasis they place on causal-informational or historical relations, as well as in the specific types of such relations they prioritise.
-Additionally, they may diverge on the inclusion of further relations, such as structural resemblance relations Shea (2018).
+In brief, internal representations are states that carry information about the world, have the selected function to do so, and are appropriately used by the system in virtue of the information they carry.
+Existing theories of basic representation vary in the emphasis they place on causal-informational relations and selection processes, as well as in the specific types of such relations they prioritise.
+Additionally, they may diverge on the inclusion of further relations, such as structural resemblance relations (Shea2018).
 Nevertheless, all of these theories, in one way or another, assert that a combination of these relations is central to the emergence of meaning in the world.
-To possess intrinsic meaning, at a minimum, is to stand in a specific set of causal-informational and historical relations with the thing in the world being represented.
+To possess intrinsic meaning, at a minimum, is to stand in a specific set of causal-informational and historical relations (via past instances of selection) with the thing in the world being represented.
 Theories of content, therefore, provide the essential ingredients that determine referential grounding, enabling internal states to hook onto things in the world.
-[^18] Philosophical discussions about cognitive representation have traditionally given considerable weight to metaphysical worries regarding the determinacy of content.
-Once seen instead as scientific, explanatory theories, we take such worries to be misguided (cf., Shea 2018, Coelho Mollo 2022).
 6 Grounding Generative Pre-trained Models
 Philosophical theories of representation identify sets of relations that imbue representations with content, giving them intrinsic meaning.
-These primarily include causal-informational and historical relations.
+These primarily include causal-informational relations and histories of selection.
 Our aim is to show that generative pre-trained models, such as LLMs, can have internal states and generate outputs with intrinsic meaning, i.e., that are referentially grounded in the world.
-[^19] For recent discussions that apply theories of representational content to LLMs, see Harding (2023), Grzankowski (2024), Goldstein & Levinstein (2024), Chalmers (2025).
+[^20] For recent discussions that apply theories of representational content to LLMs, see hardingOperationalisingRepresentationNatural2023c, grzankowskiRealSparksArtificial2024, goldsteinDoesChatGPTHave2024, chalmersPropositionalInterpretabilityArtificial2025, williams2025structuralcorrespondencesgroundreal.
 To begin, we will examine whether appropriate causal-informational relations exist between LLM representations/outputs and entities and properties in the world.
@@
 The extensive corpus of text upon which LLMs are trained bears the ‘imprint’ that the world has left on language and linguistic expression.
-A strict separation of form and meaning has been a central assumption in much of linguistics, from Ferdinand de Saussure’s principle of the “arbitrariness of the linguistic sign” (de Saussure 1916) to contemporary formal theories of syntax and semantics (Chomsky 1957, Montague 1970).
-However, this notion has faced growing criticism since the 1980s, especially in light of empirical investigation of the iconicity of language (Perniss & Vigliocco 2014).
-In this context, iconicity refers to a non-arbitrary relationship between the form of a linguistic sign and its referent, based on some similarity or analogy (Haiman 1985).
+A strict separation of form and meaning has been a central assumption in much of linguistics, from the principle of the “arbitrariness of the linguistic sign” (Saussure1916) to contemporary formal theories of syntax and semantics (chomskySyntacticStructures1957, montagueUniversalGrammar1970).
+However, this notion has faced growing criticism since the 1980s, especially in light of empirical investigation of the iconicity of language (pernissBridgeIconicityWorld2014).
+In this context, iconicity refers to a non-arbitrary relationship between the form of a linguistic sign and its referent, based on some similarity or analogy (haimanIconicitySyntax1985).
 The nature of this similarity relation can vary: imagistic iconicity involves a perceptual resemblance between a word and its referent (e.g., in onomatopoeias), while diagrammatic iconicity involves a structural resemblance.
 Diagrammatic iconicity is pervasive in linguistic structures.
-For example, the ordering of events in narrative sequences tends to reflect closeness and succession in time (e.g.
-“Veni, vidi, vici”); what is nearest to the speaker in a literal or metaphorical sense tends to be mentioned first (e.g., “here and there”); elements whose referents are closely related in some way or other tend to be adjacent in sentences (principle of adjacency); the use of subordination reflects causality or conditionality between states-of-affairs (e.g., “If you study hard, you will pass the exam”), and so on (Van Langendonck 2010).
-The pervasiveness of diagrammatic iconicity in linguistic structures suggests that the way in which linguistic signs are organised and structured is not arbitrary, and reflect some ways in which the world is organised and structured.
+For example, the ordering of events in narrative sequences tends to reflect closeness and succession in time (e.g., “Veni, vidi, vici”); what is nearest to the speaker in a literal or metaphorical sense tends to be mentioned first (e.g., “here and there”); elements whose referents are closely related in some way or other tend to be adjacent in sentences (principle of adjacency); the use of subordination reflects causality or conditionality between states of affairs (e.g., “If you study hard, you will pass the exam”), and so on (vanlangendonckIconicity2010).
+The pervasiveness of diagrammatic iconicity in linguistic structures suggests that the organization and structure of linguistic signs is not arbitrary but rather reflects some ways in which the world itself is organised and structured.
 While diagrammatic iconicity is not typically characterised in distributional terms, patterns of co-occurrence in language reflect stable statistical regularities in the way in which linguistic elements are used together.
@@
 These patterns of co-occurrence reflect the underlying structure of the world that speakers are trying to describe.
-Thus, contra Bender & Koller (2020), form and meaning are far from being so easily separable, especially at the level of general patterns of language use.
+Thus, contra Bender2020, form and meaning are far from being so easily separable, especially at the level of general patterns of language use.
 The patterns of language use derived from the training corpus reveal important aspects of the world.
 They embed information about patterns of interaction between humans and the world, human ways of categorising their environments, and the selection of features that are more or less salient to them.
-Indeed, a plausible hypothesis for LLMs’ success in a wide range of tasks is that they have induced general semantic patterns from their training data, which closely resemble what could be termed semantic ‘laws’.
-[^20] For some preliminary work on the potential form of such semantic laws, see Wolfram (2023).
-These laws are not merely formal or syntactic; rather, they tap into the intricate relationships between linguistic form and the world it describes.
-[^21] See Merrill et al.
-(2024) for concrete evidence that LLMs can induce entailment relationships between sentences from co-occurrence statistics.
+Indeed, a plausible hypothesis for LLMs’ success in a wide range of tasks is that they have induced general semantic patterns from their training data.
+These patterns are not merely formal or syntactic; rather, they tap into the intricate relationships between linguistic form and the world it describes.
+[^21] See merrillCanYouLearn2024 for concrete evidence that LLMs can induce entailment relationships between sentences from co-occurrence statistics.
 In short, LLMs are trained on data that is shaped by humans and their interactions with the world.
@@
 Many human representations are also acquired by indirect, mediated ways.
-Indeed, testimony and social learning underlie a considerable amount of the internal representations humans come to learn, and are considered to be key processes that help explain cultural knowledge and cultural evolution – in their turn central factors for the species’ success Tomasello (2009), Sterelny (2014), Heyes (2018).
+Indeed, testimony and social learning underlie a considerable amount of the internal representations humans come to acquire, and are considered to be key processes that help explain cultural knowledge and cultural evolution—in their turn central factors for the species’ success (Tomasello2009, Sterelny2014, Heyes2018).
 These include representations of entities one has not had any direct causal contact with.
-For example, most people have never been in direct causal contact with enriched uranium, but that does not entail that their thoughts and statements about uranium lack referential grounding – they are indeed about uranium.
-Representations of culturally constructed notions, such as ‘democracy’ or ‘university’, are also acquired this way – they rely on complex sets of social practices and conventions that are transmitted not by direct causal contact with documents or buildings, but rather by social learning.
+Most people have never been in direct causal contact with enriched uranium, but that does not entail that their thoughts and statements about uranium lack referential grounding—they are indeed about uranium.
+Representations of culturally constructed notions, such as ‘democracy’ or ‘university’, are also acquired this way—they rely on complex sets of social practices and conventions that are transmitted not by direct causal contact with documents or buildings, but rather by social learning.
 In a way akin to many human representations, the internal representations in LLMs are indirectly connected to the world through testimony and ‘social’ learning.
@@
 This dependence is however distinct from the one that drives the Vector Grounding Problem.
-The Problem is not how the representations of connectionist models come to acquire intrinsic meaning, but whether these representations have intrinsic meaning at all – instead of being merely ascribed meaning by external interpreters.
-To take once again the example of enriched uranium, LLMs find themselves in a position similar to most humans: their representations and outputs about enriched uranium intrinsically refer to enriched uranium – irrespective of any interpreter’s gloss – even though the process enabling such reference depends on a chain of transmission that involves other systems with intrinsic meaning.
+The Problem is not how the representations of connectionist models come to acquire intrinsic meaning, but whether these representations have intrinsic meaning at all—instead of being merely ascribed meaning by external interpreters.
+LLMs find themselves in a position similar to most humans: their representations and outputs about, say, enriched uranium intrinsically refer to enriched uranium—irrespective of any interpreter’s gloss—even though the process enabling such reference depends on a chain of transmission that involves other systems with intrinsic meaning.
 In our terminology, such internal states are non-basic representations with intrinsic meaning.
@@
 [^22] This analysis would extend, mutatis mutandis, to multimodal inputs, which are also mediated by human interactions with the world (e.g., photographs are captured by humans with a camera).
-However, one might argue that some worldly properties instantiated by visual samples in a multimodal dataset, such as shapes, can be accessed more directly than they could through the mediation of language.
-This is because images (and samples in other sensory domains) are entirely iconic, while language only exhibits limited patterns of iconicity.
 We discuss whether multimodal models fulfil the historical requirement for representational content in subsection 7.3 below.
-They are complex due to the intricate causal-informational relationships between the patterns of language use, the semantic ‘laws’ that govern them, the structure of the world they have (culturally) evolved to capture, and the causal chains of communication and testimony that produce the learning data – of which the LLM itself can be considered a link.
+They are complex due to the intricate causal-informational relationships between the patterns of language use, the structure of the world they have (culturally) evolved to capture, and the causal chains of communication and testimony that produce the learning data—of which the LLM itself can be considered a link.
 In light of these considerations, we contend that generative pre-trained models, such as LLMs, can meet the causal-informational criteria necessary for possessing referentially grounded representations and outputs.
-6.2 Historical Relations
-Although it is important to differentiate between causal-informational and historical relations when considering referential grounding, these aspects are often intertwined in real-world cases.
-The causal-informational relations previously discussed with respect to LLMs possess a historical dimension, as they depend on patterns of interaction between humans and their environments.
-Likewise, in many of the interesting cases pertaining to human cognition, the meaning-endowing causal-informational relationships can be traced back to past episodes of learning.
-That said, as discussed in section 5, finding causal-informational relations between LLMs and the world is the easiest part of our investigation, given how abundant such relations are.
-Much more challenging is to show that LLMs have internal states that not only carry information about worldly entities, but that also have the function to carry such information – so that the LLM can use it in generating appropriate outputs.
-Selection processes and learning are the historical mechanisms that theories of representation commonly invoke to explain how representations come to have the relevant function.
-These processes explain how a system acquires a specific function or purpose, in this case, the production and/or use of representations.
-Such functions clarify how cognitive systems can represent accurately – when they produce and use representations that fulfil the relevant function – as well as misrepresent, when the representations produced and/or used are inappropriate for the task.
-While it is widely accepted that Generative Pre-Trained Models undergo a learning-like process during training, it is less clear whether they undergo the relevant kind of selection process.
-It can be argued that they do, at least to the same degree as other human artefacts (Artiga 2016, Coelho Mollo 2019), given that numerous design and implementation choices form an integral part of creating artefacts such as AI systems.
-The crux of the matter, however, lies in determining whether LLMs stand in historical relations to the right sorts of things, namely the entities and properties in the world that could ground their representations Butlin (2021).
-At first glance, the situation appears analogous to that of causal-informational relations: LLMs stand in historical relations to the world, albeit in a mediated manner, through the human-generated data used for training them.
-The critical question is: can the internal states and outputs of LLMs acquire the function to represent extra-linguistic entities, despite their being trained only on language?
-In other words, the question is whether LLMs can acquire the sort of world-involving functions needed for referential grounding, and which underlie the descriptive normativity that is an essential feature of meaning.
-To address this question, it is important to distinguish between the proximate and ultimate functions performed by LLMs during training and inference.
-The proximate function of LLMs is plainly next-token prediction: predicting the next token in a text sequence, given the context provided by that sequence.
-This proximate function is infra-linguistic, and so are the standards for its fulfilment or failure, namely, whether the LLM predicts the correct tokens reliably enough.
-While such function may be suitable for underpinning relational grounding, it appears inadequate for contributing to the world-involving normativity necessary for referential grounding.
-In contrast, ultimate functions refer to the broader, more abstract goals that the model is tasked with achieving through next-token prediction.
-These functions can be world-involving in principle, such as generating text that accurately describes real-world states of affairs, or solving problems based on real-world information.
-In what follows, we will consider several potential sources for world-involving ultimate functions in the training and use of LLMs.
+6.2 Selection history
+Finding causal-informational relations between LLMs and the world is the most straightforward part of our investigation, given the abundance of such relations.
+The far greater challenge lies in showing that LLMs have internal states that not only carry information about worldly entities and properties, but also have the function of carrying such information—enabling LLMs to generate appropriate outputs based on this representational content.
+Theories of representation typically invoke selection processes, broadly construed to include learning, as the historical mechanisms that explain how internal states acquire representational functions.
+While generative pre-trained models clearly undergo a form of learning during training, whether this constitutes the appropriate kind of selection process is less clear.
+The crucial question is whether LLMs’ internal states possess the right kind of selection history—having been selected for carrying information about something and guiding behaviour accordingly—directed toward the right targets, namely the entities and properties in the world that could ground their representations (Butlin2021).
+Only when these conditions are satisfied can LLMs’ internal states acquire genuine representational functions and the descriptive normativity that accompany them.
+It is important to distinguish between the proximate and ultimate functions that LLMs perform during training and inference.
+The proximate function of LLMs is plainly next-token prediction: given a text sequence, predict the subsequent token based on contextual information.
+This proximate function is intra-linguistic, and so is its success criterion—namely, the LLM’s accuracy at predicting the correct tokens.
+While this function may support intra-linguistic relational grounding, it appears insufficient for establishing the normativity required for genuine reference to the external world.
+Ultimate functions, by contrast, encompass the broader, more abstract goals that models achieve through next-token prediction.
+These functions can, in principle, involve world-directed tasks such as generating text that accurately describes real-world states of affairs or solving problems requiring real-world knowledge.
+We will examine several potential sources of world-involving, representational ultimate functions in LLM training and deployment.
+[^23] The relevant world need not be our shared, physical world, but can encompass other extra-linguistic domains, such as the ‘game world’ of chess rules, pieces, and moves.
+When discussing this more liberal notion of ‘world-involving’, we note it explicitly in the text.
 Our general argument goes as follows.
-The training process of LLMs is a form of selection – through learning – that stabilises the functional roles of the model’s internal states.
-This process selects for internal states whose ultimate functional role extends beyond mere linguistic prediction to include extra-linguistic, world-involving functions.
+The training process of LLMs is a form of selection through learning that picks out and stabilises those internal states that lead to successful behaviour.
+In many cases, this process selects for internal states whose ultimate function extends beyond mere linguistic prediction to include extra-linguistic, world-involving correctness conditions, such as representing facts about the world.
 This is most evident in models that undergo fine-tuning to satisfy human preferences, since such fine-tuning explicitly involves selecting internal states that increase the probability of outputs satisfying world-involving norms, such as factual accuracy.
-However, we will also present evidence suggesting that pre-training alone can, in some contexts, select for internal states with world-involving content, albeit in a more indirect way.
-If LLMs indeed have internal states with extra-linguistic intrinsic meaning, and if those states causally influence the generation of linguistic outputs, then those outputs are thereby referentially grounded in the world.
+We will also investigate the possibility that, to a limited extent and within some constrained domains, pre-training on next-token prediction alone may be sufficient for endowing internal states with world-involving, representational functions.
+If LLMs indeed have internal states with extra-linguistic intrinsic meaning, and if those states are causally responsible for the generation of linguistic outputs, then those outputs are thereby referentially grounded in the world.
 This would mean that at least some outputs of LLMs are intrinsically meaningful in the sense discussed earlier: their meaning does not depend on the interpretations projected onto them by human users.
-The key premise in this argument – that the training process of LLMs can select for internal states with extra-linguistic functional roles – requires detailed defence.
-In what follows, we offer two complementary arguments supporting this premise.
-The first, which we call the argument from post-training, focuses on how preference tuning methods explicitly establish world-involving functional roles for LLM internal states.
-The second, which we call the argument from pre-training, examines how even the basic next-token prediction learning objective might, under certain conditions, implicitly select for ultimate world-involving functions.
+The key premise in this argument—that the training LLMs undergo is a kind of selection process suitable for endowing their internal states with functions to represent states of the world—requires careful defence.
+In what follows, we explore three complementary strategies to defend the premise, each aimed at showing that a specific kind of learning process in current LLMs can in principle confer world-involving, representational functions to their internal states.
+The argument from post-training focuses on fine-tuning methods based on human preferences as the most promising candidates for this role.
+The argument from pre-training investigates whether the basic next-token prediction objective might, under specific conditions, implicitly select for world-involving representational functions.
+Finally, the argument from in-context learning proposes that LLMs could, in principle, acquire world-involving representational functions during inference without parameter updates.
 6.2.1 The Argument from Post-Training
-The training process of modern LLMs, particularly those deployed in consumer-facing chatbots like ChatGPT or Claude, involves two distinct phases: pre-training and post-training.
+The training process of modern LLMs, particularly those deployed in consumer-facing chatbots like ChatGPT, involves two distinct phases: pre-training and post-training.
 Pre-training involves training the base model from scratch on a next-token prediction objective across vast corpora of text.
 By contrast, post-training adapts this base model to behave in ways that go beyond generating merely plausible text completions, making LLMs more useful for various downstream tasks and applications.
-The post-training phase typically includes two types of fine-tuning: instruction tuning, and preference tuning.
-Instruction tuning is a form of supervised fine-tuning on curated instruction-response pairs (Ouyang et al.
-2022).
+This phase typically includes two types of fine-tuning: instruction tuning and preference tuning.
+Instruction tuning is a form of supervised fine-tuning on curated instruction-response pairs (ouyangTrainingLanguageModels2022).
 Given an instruction, the model is trained to generate the desired output by minimising the difference between its predictions and the target responses in the instruction dataset.
 This teaches the model to generate appropriate responses that follow user instructions rather than simply generating plausible completions of text sequences.
-Preference tuning, by contrast, trains LLMs to generate outputs that better align with selected normative criteria, such as helpfulness, harmlessness, and honesty or factual accuracy (Askell et al.
-2021).
+Preference tuning further trains LLMs to generate outputs that better align with selected normative criteria, such as helpfulness, harmlessness, and honesty or factual accuracy (askellGeneralLanguageAssistant2021).
 Unlike standard supervised fine-tuning, which only demonstrates what to do, preference tuning explicitly trains models to distinguish between better and worse responses according to these criteria.
 This is typically implemented by training on pairs of responses where human evaluators have indicated a preference for one over the other.
-Reinforcement Learning from Human Feedback (RLHF) is the most well-known preference tuning method (Christiano et al.
-2017).
+Reinforcement Learning from Human Feedback (RLHF) is the most well-known preference tuning method (christianoDeepReinforcementLearning2017).
 RLHF begins by collecting human preferences over model outputs for a diverse set of prompts.
 These preferences are used to train a reward model that predicts how humans would evaluate various responses.
-The LLM is then fine-tuned using reinforcement learning to maximize this reward function while remaining reasonably close to its original behaviour.
-Another increasingly popular approach is Direct Preference Optimization (DPO), which achieves similar results without explicitly modelling the reward function (Rafailov et al.
-2024).
+The LLM is then fine-tuned using reinforcement learning to maximise this reward function while remaining reasonably close to its original behaviour.
+Another increasingly popular approach is Direct Preference optimisation (DPO), which achieves similar results without explicitly modelling the reward function (rafailovDirectPreferenceOptimization2024).
 Instead, DPO directly optimises the model to assign higher probabilities to preferred responses and lower probabilities to rejected ones.
-When an LLM is optimised according to human preferences that incorporate extra-linguistic normative standards like factual accuracy, this creates a selection process that reinforces specific activation patterns based on their relationship to worldly states.
-These internal states acquire a specific functional role: to track relevant aspects of the world that enable the model to generate outputs that satisfy human preferences for truthfulness.
-To illustrate how preference tuning establishes world-involving functions, consider a toy example focused solely on factual accuracy using DPO.
+When an LLM is optimised according to human preferences, a selection process is established that reinforces specific internal states based on their contribution to externally rewarded outputs.
+Consider a toy example using DPO.
 We begin with a language model that has undergone both pre-training and instruction tuning.
@@
 When we apply DPO to this model, it adjusts the model’s parameters to increase the probability of generating factually accurate responses and decrease the probability of generating less accurate ones.
-[^23] Note that DPO also incorporates a constraint that prevents the model from deviating too far from its original behaviour, to maintain the right balance between preference satisfaction and fluency.
-Through this process, the model’s internal states are selected specifically because they enable the model to satisfy the norm of factual accuracy, thereby providing a standard of success that depends not on linguistic patterns alone, but on the correspondence between the generated text and states of affairs in the world.
-This functional role provides another piece of the story – in addition to causal-informational relations – of how LLM internal states can be referentially grounded, and thus overcome the Vector Grounding Problem.
-The internal states of the preference-tuned LLM represent aspects of the world because they were selected specifically for their ability to carry information about those aspects in service of generating truthful outputs.
-When the model succeeds in generating a factually accurate response, its internal representations are fulfilling their function of tracking relevant worldly states.
-Importantly, the accuracy of the model is not what determines whether its internal states are referentially grounded.
-Even a partially inaccurate model prone to occasional hallucinations can have internal states with world-involving content, provided those states were selected for their ability to carry information about aspects of the world.
-What matters for grounding is not perfect accuracy, but rather the presence of a representational function with world-involving correctness conditions.
-Preference tuning thus provides a clear mechanism by which LLMs can acquire internal states with extra-linguistic functional roles, insofar as RLHF or DPO incorporate factual accuracy as an explicit criterion over rewarded preferences.
-[^24] At least in those domains where such factual accuracy has been emphasized during preference tuning.
-While we focus on factual accuracy, all three common norms in preference tuning – the “3 Hs” of helpfulness, harmlessness, and honesty – are arguably world-involving.
-Helpfulness requires generating outputs that genuinely assist users with their goals, which depends on real-world task success conditions rather than mere linguistic coherence.
+Through this process, the model’s internal states are selected because they enable the model to satisfy the norm of factual accuracy, thereby providing a standard of success that depends not on linguistic patterns alone, but on the correspondence between the generated text and states of affairs in the world.
+This provides the missing piece of the story—in addition to causal-informational relations—of how the internal states of LLMs can be referentially grounded, and thus overcome the Vector Grounding Problem.
+The internal states of the preference-tuned LLM represent states of the world because they were selected specifically for their ability to carry information about those states in service of generating truthful outputs.
+When the model succeeds in generating a factually accurate response, its internal states fulfil their function of representing relevant worldly states.
+Two clarifications are in order.
+Firstly, we should distinguish between the function of the learning mechanism and the function of the representational states selected by that mechanism.
+The designers’ intentions bestow an artefactual function on the learning mechanism (e.g., the preference tuning pipeline has the intended function of updating parameters based on reward signals).
+However, the function of the model’s internal states is set by the learning process itself.
+The situation is analogous to artificial selection in biology.
+A dog breeder’s intention to produce calmer dogs establishes a selection regime, but it is the objective history of differential reproduction that confers the biological function of producing calmness on the selected traits.
+This function is a fact about the lineage’s history, not the breeder’s mind.
+Similarly, the designers of an LLM create a selection apparatus, but the autonomous operation of this apparatus on data confers new, selected-effects functions on the model’s internal states.
+The normativity—the standard for correctness or incorrectness—is supplied by this objective history of selection, not the designers’ intentions.
+Secondly, the extra-linguistic content of the model’s representations is determined by the specific worldly conditions that drove their selection, which in the case of preference tuning means whatever properties the reward signal actually and reliably tracks.
+In that regard, the focus on factual accuracy in our example is a special case.
+When human evaluators consistently provide higher rewards for factually accurate statements, the selection process picks out and stabilises the model’s internal states that carry relevant factual information about the world, in virtue of their doing so.
+The other two common norms in preference tuning—helpfulness and harmlessness—are also suitable to selecting referentially grounded internal states.
+Helpfulness requires generating outputs that genuinely assist users with their goals, which hinges on real-world task success conditions rather than mere linguistic coherence.
 Harmlessness involves avoiding outputs that could lead to real-world negative consequences, requiring sensitivity to causal relationships between linguistic outputs and potential worldly outcomes.
-Even instruction tuning alone could potentially establish world-involving functions, as successfully following instructions often requires tracking aspects of the world that would make the instruction completion successful.
+Even instruction tuning alone could establish world-involving functions, as successfully following instructions often requires tracking aspects of the world that would make the instruction completion successful.
 Many instructions implicitly contain success conditions that depend on extra-linguistic factors.
-Be that as it may, factual accuracy provides the most direct and straightforward case of world-involving content, where the connection between internal states and worldly conditions is the most evident.
-It offers the clearest case for how LLM internal states can be referentially grounded, and thus endowed with intrinsic meaning.
+In summary, post-training establishes the selection processes necessary for endowing LLM internal states with representational functions, which together with causal-informational relations are the key ingredients for intrinsic meaning.
+It is worth noting that the specific norms used in post-training shape what the internal states represent.
+Achieving referential grounding does not automatically guarantee that the meanings of the internal representations and outputs of LLMs align with their conventional human meanings.
+Un-humanlike selection criteria, together with the other un-humanlike features of LLMs, may endow them with ‘deviant meanings’.
+This is an intriguing possibility that lies beyond the scope of this paper.
+What matters for our purposes is that LLMs’ internal states and outputs can be intrinsically meaningful, even if such meanings may be alien to humans.
 6.2.2 The Argument from Pre-Training
-We now turn to a more controversial claim: that pre-training alone can, in certain contexts, establish world-involving functions.
-During pre-training, LLMs are explicitly trained to minimise the negative log-likelihood of the next token given previous tokens – an objective limited to the statistical distribution of tokens in the training corpus.
-[^25] See McCoy et al.
-(2024) for an approach to studying LLMs, which they call the “the teleological approach”, that takes as its starting point this proximate function.
+We now turn to a more controversial claim: that pre-training alone can, under specific conditions, establish world-involving representational functions.
+The proximate function of an LLM during pre-training is to minimise the negative log-likelihood of the next token given a preceding sequence.
+[^24] See McCoy2024 for an approach to studying LLMs, which they call the “the teleological approach”, that takes as its starting point this proximate function.
 They show that LLM behaviour is strongly shaped by the distributional statistics of the training dataset.
@@
 Sensitivity to distributional patterns can be a means to fulfilling ultimate goals, given all the worldly information embedded in such patterns.
-At first glance, this appears to be a purely infra-linguistic task with no connection to the world.
-However, as we have seen in subsection 6.1, the patterns of language use in the training corpus are not arbitrary; they reflect systematic regularities in how humans use language to communicate about the world.
-From an information-theoretic perspective, there are vastly more ways to linguistically express world states than there are world states themselves.
-For next-token prediction, two strategies are possible: direct memorisation of observed sequences, or latent variable factorisation that learns to represent underlying causal factors.
-The parameter efficiency of the latter strategy relative to the former scales with the ratio of linguistic expressions to world states – typically a very large number.
-This creates a strong selection pressure for internal states that track latent variables, as they enable more efficient prediction across diverse linguistic contexts.
-The Information Bottleneck principle provides a formal framework for understanding this process (Tishby & Zaslavsky 2015).
-Neural networks learn to extract approximate, minimally sufficient statistics of the input with respect to the output.
-This involves finding a maximally compressed mapping of the input that preserves as much information as possible about the output.
-In language models, this connection is even more fundamental.
-As shown by Delétang et al.
-(2024), minimising the next-token prediction loss is mathematically equivalent to finding the most efficient way to compress tokenised information.
-For complex data distributions generated by structured world domains, the most efficient compression involves factorising the problem according to the causal structure that generated the data.
-Research in mechanistic interpretability provides concrete evidence that models pre-trained on next-token prediction can develop internal representations with extra-linguistic content at least in some domains.
-One particularly interesting case is Othello-GPT, a transformer model trained solely to predict moves in Othello game transcripts (Li et al.
-2023).
+At first glance, this appears to be a purely intra-linguistic objective, with correctness conditions defined entirely by the statistical distribution of tokens in the training corpus.
+It is not immediately obvious how a selection process governed by such a criterion could give rise to internal states whose ultimate function is to represent extra-linguistic reality.
+To succeed at next-token prediction, two strategies are possible: memorising observed sequences or inducing latent factors responsible for those sequences.
+For large datasets, the latter strategy is far more efficient (deletangLanguageModelingCompression2024).
+This can create a selection pressure for internal states that track underlying causal factors, as they enable more efficient prediction across diverse linguistic contexts.
+Such selection pressure may be enough to endow language models with internal states that stand in mapping relations to states of affairs in the world, sometimes called ‘world models’ (yildirimTaskStructuresWorld2023).
+However, philosophers of representation have long recognised that standing in mapping relations to states of the world—also known as ‘structural resemblance’—is insufficient for representation.
+Such mappings are cheap: they can be found between any two sets with the same cardinality.
+Representational content would be trivial and ubiquitous if structural resemblance were sufficient to secure it (Shea2012-SHEMIR).
+For internal states that merely map onto extra-linguistic domains to become full-fledged representations, they must have the function of representing those worldly domains and be used by the system to generate outputs precisely because they have this representational function (Shea2018, williams2025structuralcorrespondencesgroundreal).
+The concept of meta-learning provides a formal framework for understanding how pre-training could, in principle, establish world-involving representational functions.
+Meta‑learning, or ‘learning to learn’, is a bi‑level optimisation process with two learning loops: the outer loop trains a model over a range of related tasks, while the inner loop refers to the model’s rapid adaptation when it encounters data from a specific task.
+Rather than seeking perfect performance on any one training task, the outer loop shapes the model’s parameters so that a small number of gradient steps (the inner loop) can quickly tailor those parameters to a new, previously unseen task from the same distribution.
+The pre-training of LLMs can be viewed as an implicit form of meta-learning (wuMetaLearningPerspectiveTransformers2024).
+The outer loop is the standard training process: minimising next-token prediction loss via gradient descent on the model’s parameters over a large corpus.
+This corpus implicitly contains a massive distribution of distinct tasks—translating, answering questions, summarising, solving problems, playing games, etc.
+The outer loop thus selects for a single set of parameters that makes the model an effective learner for the myriad tasks embedded in its training data.
+Wu et al.
+suggest that the inner loop is the forward pass itself.
+On their analysis, the fixed parameters learned during the outer loop do not encode a static knowledge base, but rather an on-the-fly optimisation algorithm.
+For any given input sequence (i.e., a task context), the forward pass through the Transformer’s layers functions as an iterative optimisation process.
+Each layer takes the token representations from the layer below and performs an update, which Wu et al.
+show can be mathematically interpreted as an optimisation step that refines the representations to make them progressively more suitable for the final next-token prediction.
+This perspective provides a plausible, though speculative, mechanism for the emergence of world-involving representational functions during pre-training.
+The outer loop’s general, domain-agnostic objective (predict the next token) exerts a selection pressure on the dynamics of the inner-loop optimiser (the forward pass).
+For specific domains represented in the training data, the most efficient strategy for the inner loop to succeed at its task-specific prediction objective may be to construct and manipulate an internal model of the latent, extra-linguistic state of affairs that generated the data.
+The proximate, intra-linguistic objective of the outer loop would thereby select for an inner-loop process whose success is conditional on tracking extra-linguistic states.
+In such cases, the resulting internal states would acquire the ultimate function of representing that extra-linguistic domain.
+Their function would be to carry information about the world because their having been selected by the learning process is explained by the fact that they did so, thereby enabling more accurate predictions.
+The argument’s viability rests on two substantive assumptions.
+The first is that the meta-learning perspective provides a valid mechanistic account of the Transformer’s forward pass during pre-training.
+The second is that for certain domains, the general pre-training objective (next-token prediction) creates selection pressures for an inner-loop process whose success is conditional upon tracking extra-linguistic objectives.
+The argument is most plausible for what we might call formally constrained domains, such as logic puzzles or board games.
+In such domains, the condition specified in the second assumption is most clearly met.
+The sequences of tokens in the training data (e.g., game moves) are not generated by the contingent communicative intentions of humans in an open-ended world, but are instead governed by objective, formal rules that constitute the domain itself.
+The statistical regularities in the data are direct consequences of this underlying formal structure.
+For a model to succeed at its proximate, intra-linguistic task of predicting valid next tokens, it must become sensitive to these non-accidental regularities.
+The most robust and efficient strategy for achieving this is to induce an internal model of the domain’s rules and states—the very structure that generates the data.
+This establishes a non-arbitrary, world-involving standard of correctness, allowing the inner-loop selection process to endow internal states with the ultimate function of representing that extra-linguistic domain.
+Research in mechanistic interpretability provides concrete evidence that models pre-trained on next-token prediction can develop internal representations with extra-linguistic content at least in formally constrained domains.
+One particularly interesting case is Othello-GPT, a transformer model trained solely to predict moves in Othello game transcripts (liEmergentWorldRepresentations2023).
 Despite receiving no explicit information about Othello’s rules or board structure, this model was found to develop internal representations that encode the current game state in a player-centric manner.
-Nanda et al.
-(2023) showed through intervention studies that specific directions in the model’s activation space encode board positions as "Mine," "Yours," or "Empty." When they intervened on these representations by ’flipping’ a tile from ’yours’ to ’mine’ in the model’s activation space, the model’s predictions changed in ways consistent with the altered board state.
+nandaEmergentLinearRepresentations2023 showed through intervention studies that specific directions in the model’s activation space encode board positions as ‘Mine’, ‘Yours’, or ‘Empty’.
+When they intervened on these representations by ‘flipping’ a tile from ‘yours’ to ‘mine’ in the model’s activation space, the model’s predictions changed in ways consistent with the altered board state.
 This suggests not only that the model encodes the board state, but that this encoding plays a causal role in determining its predictions of legal next moves.
-[^26] For a longer philosophical discussion of causal interventions used to understand interrnal states of Othello-GPT, see Millière & Buckner (forthcoming).
-The Othello-GPT example illustrates how selection processes during pre-training can establish ultimate functional roles with extra-linguistic correctness conditions.
+[^25] For a longer philosophical discussion of causal interventions used to understand internal states of Othello-GPT, see milliereInterventionistMethodsInterpreting2025.
+The Othello-GPT example illustrates how selection processes during pre-training may establish representational functions with extra-linguistic correctness conditions.
 During training, the model developed specific directions in its activation space that linearly encode the board state from the perspective of the current player.
-These linear representations emerged because they enabled the model to accurately predict legal moves across diverse game contexts.
-Their function became to carry information about the current state of the Othello board, with accuracy conditions dependent on the actual game state rather than merely on linguistic patterns.
-This example is particularly compelling because game transcripts are systematically constrained by the rules of Othello – to predict legal moves effectively, the model developed a linearly decodable internal model that tracks the underlying game state.
+These linear representations plausibly emerged and persisted because they enabled the model to accurately predict legal moves across diverse game contexts.
+Their function thus became that of carrying information about the current state of the Othello board, with accuracy conditions dependent on the actual game state and the learned rules, rather than merely on linguistic patterns.
+This example is particularly compelling because game transcripts are systematically constrained by the rules of Othello—to predict legal moves effectively, the model developed a linearly decodable internal model that tracks the underlying game state.
+[^26] Further research suggests that Othello-GPT may learn many independent, localised decision rules (a “bag of heuristics”) rather than a single systematic algorithm for tracking the board state and determining legal moves (jylin04OthelloGPTLearnedBag2024).
+For instance, researchers identified specific “board pattern neurons” that activate when a particular local arrangement of pieces makes a specific move legal.
+This is consistent with our argument; the crucial point is that the internal states guiding such heuristics acquire a world-involving representational function through their selection history.
+A given “board pattern neuron” was selected by the training process because its activation reliably correlated with a specific, rule-governed state of affairs on the Othello board, and detecting this state of affairs improved the model’s ability to predict a legal next move.
+The correctness condition for this neuron’s firing is therefore extra-linguistic: it is supplied by the actual state of the Othello board as governed by the game’s objective rules.
+If the neuron fires when the pattern is not present, or fails to fire when it is, it has malfunctioned with respect to the game of Othello, and this malfunction increases the probability of being penalised by the loss function.
 The case of LLMs trained on diverse corpora is undoubtedly more complex than Othello-GPT.
@@
 By contrast, general LLMs encounter messy, open-ended data with few rigid underlying rules.
-To the extent that language models may acquire internal states with world-involving content, similarly to Othello-GPT with respect to the game world, this presumably only happens in limited domains and contexts, in which next-token prediction on text sequences provides the right kind of learning signal.
-That said, there is converging evidence from mechanistic interpretability research that LLMs have internal representations that carry information about discourse entities and their relations, and that such representations are causally relevant to the generation of outputs (e.g., Li et al.
-2021; see Millière & Buckner forthcoming for a review).
-A useful perspective for understanding how pre-training can establish world-involving functional roles is to view it through the lens of meta-learning.
-Meta-learning – often characterized as ‘learning to learn’ – is the process by which a model improves its learning capabilities through experience across multiple tasks, essentially learning how to learn more efficiently.
-The pre-training of LLMs can be viewed as a form of meta-learning because it trains the model on a diverse distribution of tasks implicitly contained within the next-token prediction objective.
-When predicting the next token in various contexts across a large corpus, the model encounters many implicit ‘tasks’, ranging from syntactic completion to factual recall, reasoning, and various domain-specific problems.
-This defines a bi-level optimisation structure characteristic of meta-learning: an outer loop (the training process itself) that optimises model parameters, and an inner loop (the forward pass during inference) that applies these learned parameters to solve new tasks on the fly (Wu & Varshney 2024).
-This helps explain why pre-trained LLMs can rapidly adapt to new tasks from just a few examples at inference time, a capacity known as in-context learning (ICL): they have essentially learned optimisation strategies that generalise across the diverse task distribution encountered during pre-training, enabling efficient transfer to novel contexts during inference.
-Many different approaches to understanding in-context learning have been proposed.
-One prominent theory suggests that ICL involves implicitly optimising over latent task objectives.
-For example, von Oswald et al.
-(2024) showed that autoregressive transformers trained on sequence prediction tasks develop gradient-based optimisation algorithms in their forward pass.
-This suggests that when pre-trained LLMs encounter a series of examples in their context window, they can perform an implicit optimisation process over the latent task space.
-In other words, they can identify the underlying task structure from the examples, and construct an internal objective that captures the task function.
-When the in-context examples reflect an extra-linguistic task, the implicit objective function defined by the task may also be extra-linguistic.
-[^27] See, for example, Patel & Pavlick (2022)’s experiment which involves prompting the model in a few-short learning regime to learn mappings from colour terms like ‘cyan’ to points in RGB colour space such as [0, 100, 100].
-This is an instance of a world-involving function, having a standard of success and failure directly contingent on how the world is.
-Optimising this implicit function during the forward pass should dynamically select activation patterns that increase the likelihood of outputs satisfying the world-involving task constraints.
-In summary, the argument from pre-training suggests that even without post-training methods like preference tuning, language models may develop internal representations with ultimate world-involving functions through the selective process of next-token prediction over structured text data, at least in some limited domains.
-This can occur because inducing the latent causal structure of the data provides an informationally efficient solution to the prediction task across diverse contexts.
-Evidence from mechanistic interpretability studies supports this view at least in toy domains, showing that Generative Pre-Trained Models can develop internal states that carry information about extra-linguistic properties and causally influence outputs.
-Furthermore, the meta-learning perspective on in-context learning suggests that pre-trained LLMs can implicitly optimise ultimate extra-linguistic task functions at inference time without parameter update.
+To the extent that language models may acquire internal states with world-involving content purely through pre-training, similarly to Othello-GPT with respect to the game world, this presumably only happens in formally constrained domains, in which next-token prediction on text sequences provides the right kind of learning signal—provided that the two aforementioned assumptions hold.
+In summary, the argument from pre-training suggests that even without post-training methods like preference tuning, language models may develop internal representations with ultimate extra-linguistic representational functions through the selective process of next-token prediction over structured text data, at least in some limited, formally constrained domains.
+Evidence from mechanistic interpretability studies supports this view at least in simple game domains, showing that generative pre-trained models can develop internal states that carry information about extra-linguistic properties and causally influence outputs.
+6.2.3 The Argument from In-Context Learning
+After pre-training, LLMs can rapidly adapt to new tasks from just a few examples at inference time, a capacity known as in-context learning (ICL).
+Through ICL, pre-trained LLMs appear capable of ‘learning’ various tasks on the fly without any update to their internal parameters.
+These tasks can be intra-linguistic (e.g., translating between languages) or extra-linguistic (e.g., answering questions about world capitals).
+An intriguing question is whether ICL can be sufficient to achieve referential grounding.
+Suppose that we invent a new board game called ‘Chroma’, played by placing coloured tiles on a grid.
+The game has a clear set of rules: for instance, a blue tile can only be placed adjacent to a red or yellow tile, and scoring depends on creating specific geometric patterns.
+These rules define the game’s ‘world’—a self-contained domain with its own entities (tiles, grid positions), properties (colours), and causal regularities (legal moves, scoring conditions).
+Now, imagine we present a pre-trained LLM, which has never encountered Chroma before, with a prompt containing the complete rules of the game, followed by a transcript of the first few moves.
+We then prompt it to generate the next legal and strategically optimal move.
+For the LLM to succeed at this task, it cannot rely on merely matching statistical patterns from its pre-training data, as Chroma is a novel game.
+It must, in some sense, induce and apply the rules of the game without further training.
+To the extent that an LLM can do this through ICL, we argue that this can be understood as a selection process that establishes world-involving representational functions for the model’s internal representations.
+Recent research on ICL suggests that in formal domains, Transformers can learn to use the in-context examples to construct an implicit, task-specific objective function and then perform computations functionally equivalent to gradient descent to find a solution—a process called ‘mesa-optimisation’ (oswaldUncoveringMesaoptimizationAlgorithms2024).
+[^27] Mesa-optimisation at inference time is analogous to the aforementioned inner-loop optimisation during pre-training.
+The initial layers of the model parse the context and construct what is effectively a training set within the model’s activation space.
+Subsequent layers then iteratively refine the internal representations, taking steps to minimise an implicit loss function tailored to the task at hand.
+During ICL, this mesa-optimisation process within the forward pass constitutes a selection process that provides a transient locus of normativity for the system’s internal states, even without parameter updates.
+It selects for internal activation patterns because they contribute to satisfying the task function defined in context.
+In the Chroma example, this task function is plausibly extra-linguistic with respect to the game-world.
+The objective that would have to be mesa-optimised is not to predict a linguistically plausible next token, but to identify a move that is valid according to the game rules.
+In other words, the correctness conditions for this task are determined not by the statistics of English, but by the objective state of the Chroma game world.
+For example, an activation pattern that correctly represents that “placing a blue tile at C4 is a legal move” will be reinforced and stabilised during the forward pass because it helps to minimise the implicit error against the objective of making a valid move.
+Conversely, a pattern representing an illegal move will be suppressed.
+These internal states thereby acquire a transient selected-effects function: their function is to carry information about the state of the Chroma game (e.g., the board configuration, legal moves, strategic opportunities) in service of winning the game.
+This function, and the normativity it entails, is established on-the-fly, for the duration of solving the specific task presented in context, and will cease to exist once that context is deleted or overwritten.
+Whether this actually occurs in LLMs is an open question.
+The emergence of effective mesa-optimisers is not guaranteed and depends on the properties of the pre-training data distribution (zhengMesaOptimizationAutoregressivelyTrained2024).
+Nonetheless, the argument from in-context learning provides a speculative but mechanistically plausible route by which an LLM could, in principle, acquire short-lived representations with extra-linguistic content on the fly.
 7 Implications
 7.1 Parameter-identical duplicates
-The historical component of our account of referential grounding invites consideration of a classic philosophical objection.
+The historical component of referential grounding invites consideration of a classic philosophical objection.
 The ‘swampman’ thought experiment envisions a perfect duplicate of a human spontaneously formed by random chance (in the original thought experiment, by a lightning strike on a swamp), and thus without any causal history.
@@
 First, the internal states of the swamp system, at least at the moment of creation, would lack any causal-informational relations to the world, given that, by hypothesis, the system has no causal history.
-Moreover, the absence of a selection and learning history means that the internal states of the swamp system lack any functions, including functions to carry information about the world Shea (2018).
-Consequently, there is no principled way to determine which outcomes count as successes versus failures for a swamp system.
-$M_{2}$ might produce outputs that appear meaningful to human interpreters, but without the selective history that shaped $M_{1}$, there is no objective basis for treating some of its outputs as ‘correct’ and others as ‘incorrect’.
-The parameter values themselves, divorced from the process that selected them, cannot ground representational content.
-[^28] This scenario is somewhat akin to the example, mentioned above, of ants forming the shape of words in the sand by pure chance; or of chimps eventually reproducing a Shakespeare play by randomly hitting keys in a typewriter.
-Such outputs have no intrinsic meaning, as they were not produced in the appropriate ways, though referentially grounded beings, such as ourselves, may extrinsically impose meaning on them.
+Moreover, the absence of a selection and learning history means that the internal states of the swamp system lack any functions, including functions to carry information about the world (Shea2018).
+Consequently, there is no principled way to determine which outcomes count as successes or failures for a swamp system.
+$M_{2}$ might produce outputs that appear meaningful to human interpreters, but without the selection history that shaped $M_{1}$, there is no objective basis for treating some of its outputs as ‘correct’ and others as ‘incorrect’.
+The parameter values themselves cannot ground representational content.
 It is worth noting, however, that referential grounding could be established relatively quickly if $M_{2}$ were subsequently fine-tuned.
-Even a brief learning history would begin to establish the relevant world-involving functions Shea (2018).
+Even a brief learning history would begin to establish the relevant world-involving functions (Shea2018).
 What matters for grounding is not the parameter values per se, but the historical process through which those parameters were selected to perform specific functions.
-We can also consider a variant of this thought experiment: [^29] We thank an anonymous referee for raising this scenario.
+[^28] A subtle variant of this case is a Swamp LLM with frozen weights that would nonetheless engage in in-context learning, adapting its outputs based on information provided within a conversational context.
+One might be tempted to argue that this ‘inner loop’ of adaptation within the forward pass constitutes a sufficient selection process to establish world-involving functions ab initio.
+We contend that it does not.
+The capacity for ICL is itself a complex functional property that requires a selection history.
+In a normally trained model, this capacity is the product of the ‘outer loop’ of pre-training, which selects for internal mechanisms that can efficiently adapt to the diverse tasks and statistical patterns encountered in the training data.
+The ICL mechanism in such a model thus has a selected-effects function: to enable rapid task adaptation based on contextual cues.
+A Swamp LLM, by contrast, possesses this mechanism by chance, so the mechanism itself lacks any function.
+Its apparent adaptation is a form of pseudo-learning—a causal process that mimics genuine, normatively-grounded adaptation but lacks the historical basis for correctness conditions.
+Without a function to represent correctly, its internal states cannot misrepresent, and therefore cannot be said to represent at all.
+We are grateful to Harvey Lederman for suggesting this example.
+We can also consider a variant of this thought experiment:
 Lucky Pre-Training.
@@
 $M_{2}$ undergoes only pre-training on next-token prediction, but through some extraordinary coincidence, ends up with exactly the same parameter values as the fully trained $M_{1}$.
-This scenario differs importantly from the standard swampman case, as $M_{2}$ does have a learning history – it just differs from that of $M_{1}$.
+This scenario differs importantly from the standard swampman case, as $M_{2}$ does have a learning history—it just differs from that of $M_{1}$.
 The answer to whether $M_{2}$ achieves referential grounding depends on whether one accepts our argument from pre-training in Section 6.2.2.
-If one accepts that pre-training alone can, in some cases, establish world-involving functional roles (as we argued), then $M_{2}$ could achieve referential grounding despite skipping the post-training phase.
-However, the fact that $M_{2}$ happens to have identical parameter values to $M_{1}$ would be a mere coincidence irrelevant to the grounding question.
+If one accepts that pre-training alone can, in some cases, establish world-involving representational functions, then $M_{2}$ could achieve referential grounding despite skipping the post-training phase.
+The fact that $M_{2}$ happens to have identical parameter values to $M_{1}$ would be a mere coincidence irrelevant to the grounding question.
 Indeed, its referential grounding would be limited to that achieved via pre-training, while the internal states that drew their grounding from preference tuning in $M_{1}$ would be ungrounded in $M_{2}$.
-These thought experiments highlight the crucial role that historical relations play in referential grounding.
+These thought experiments highlight the crucial role that a history of selection plays in referential grounding.
 The internal states of a system, no matter how complex or functionally organised, cannot have representational content without the appropriate history.
-At the same time, our position avoids an overly restrictive view of what learning histories can establish referential grounding, allowing that different training regimes might enable a system to overcome the Vector Grounding Problem through different causal-historical paths.
+At the same time, our position avoids an overly restrictive view of which learning histories can establish referential grounding, allowing that different training regimes might enable a system to overcome the Vector Grounding Problem through different causal-historical paths.
 7.2 The impact of fine-tuning
-In a recent paper, Grindrod (2024) raises an objection to our argument from post-training.
-Grindrod acknowledges our claim that fine-tuning with human feedback introduces world-involving functions that could in principle establish the necessary historical relations for referential grounding.
+In a recent paper, grindrodLargeLanguageModels2024 raises an objection to our argument from post-training.
+Grindrod acknowledges our claim that fine-tuning with human feedback introduces selection processes that could in principle establish the necessary historical relations for referential grounding.
 However, he objects:
-The problem with this idea is that it would be odd to claim that a relatively marginal procedure that is designed to optimize the performance of a LLM on specific tasks ends up being the difference between intentionality and its lack.
-The reason, simply put, is that the fine-tuning procedures – whether RLHF, supervised training, or something else – are not the reason why we have such impressive LLMs today.
-(Grindrod 2024, p.
+The problem with this idea is that it would be odd to claim that a relatively marginal procedure that is designed to optimise the performance of a LLM on specific tasks ends up being the difference between intentionality and its lack.
+The reason, simply put, is that the fine-tuning procedures—whether RLHF, supervised training, or something else—are not the reason why we have such impressive LLMs today.
+(grindrodLargeLanguageModels2024, p.
 71)
@@
 We are not claiming that preference tuning explains the impressive capabilities of LLMs or their ability to produce contextually appropriate outputs.
-Rather, we are specifically addressing how LLMs can overcome the Vector Grounding Problem – how their internal states can possess intrinsic meaning by being referentially grounded in the world.
+Rather, we are specifically addressing how LLMs can overcome the Vector Grounding Problem—how their internal states can possess intrinsic meaning by being referentially grounded in the world.
 These are distinct questions.
@@
 The quality of outputs and their status as intrinsically meaningful are orthogonal issues.
-What matters for referential grounding is not how much a particular training procedure contributes to performance quality, but whether it establishes the right kind of functional relations between internal states and worldly entities.
+What matters for referential grounding is not how much a particular training procedure contributes to performance quality, but whether it helps to establish the right kinds of relations between internal states and worldly entities.
 Even if preference tuning procedures make only marginal contributions to an LLM’s overall capabilities, they can fundamentally alter the normative conditions under which the model operates by introducing world-involving correctness standards.
 7.3 Multimodality
-The preceding discussion has identified two potential sources of referential grounding for LLMs: (1) preference tuning with human feedback that incorporates world-involving criteria such as factual accuracy, and (2) pre-training on next-token prediction, which may select for internal states with world-involving functional roles in limited domains where this is the most efficient way to achieve good prediction accuracy.
-When we turn to multimodal and embodied systems, we need to examine whether and how these sources of referential grounding extend to such systems.
-Our analysis leads to a surprising conclusion: multimodality and embodiment are neither necessary nor sufficient conditions for referential grounding in artificial systems, though they can complement and potentially enhance referential grounding when the right kind of learning process is available.
-Let us first consider a simple multimodal model like CLIP (Radford et al.
-2021).
+At first blush, multimodal and embodied AI systems seem to be the most obvious candidates for overcoming the Vector Grounding Problem.
+Unlike pure LLMs, they receive sensory inputs (e.g., pixels) and/or interact physically with the world, so that the sort of ‘sensorimotor contact’ we discussed in subsection 4.2 seems to obtain.
+In section 5, we pointed out that an account of grounding based on mere causal contact is incomplete.
+Indeed, it follows from our analysis in section 6 that multimodality and embodiment are neither necessary nor sufficient for referential grounding, though they can be part of what makes referential grounding obtain when combined with the right kind of learning process.
+Let us first consider a simple multimodal model like CLIP (radfordLearningTransferableVisual2021).
 CLIP is developed by jointly training an image encoder and a text encoder to predict the correct pairings of images and text captions using a contrastive learning objective.
-During training, CLIP receives batches of image-text pairs and learns to maximize the similarity between embeddings of matched pairs while minimizing the similarity between embeddings of unmatched pairs.
-This training approach allows CLIP to develop representations that capture semantic relationships between visual and linguistic content.
-Through contrastive pre-training with image-caption pairs, CLIP achieves what we called sensorimotor grounding, albeit to a very limited extent compared to biological organisms – namely connecting linguistic vector representations to pictorial vector representations.
-CLIP also arguably achieves some degree of relational grounding in the purely linguistic domain, although there is evidence that contrastive learning from image-caption pairs severely limits the degree to which this kind of model can induce syntactic structure compared to text-only LLMs (Yuksekgonul et al.
-2023).
+During training, CLIP receives batches of image-text pairs and learns to maximise the similarity between embeddings of matched pairs while minimising the similarity between embeddings of unmatched pairs.
 As far as causal-informational relations go, the account we gave for LLMs should also apply to models like CLIP: since both images and their captions are shaped by human interactions with the world, CLIP should stand in complex, mediated causal-informational relations to the world.
 The images in CLIP’s training data were produced by humans using cameras to capture real-world scenes, and the captions were written by humans to describe those scenes.
-This establishes an indirect causal chain connecting CLIP’s internal representations to worldly entities and properties.
-However, the contrastive learning objective optimises purely for matching images with their captions, without establishing clear world-involving functional roles for CLIP’s internal states.
+This establishes an indirect causal chain connecting CLIP’s internal states to worldly entities and properties.
+However, the contrastive learning objective optimises purely for matching images with their captions, without establishing world-involving functions for CLIP’s internal states.
 There is no obvious selection pressure for internal states that track features of the world beyond what is necessary for the matching task, and no clear normative standard for success beyond the matching task itself.
-Thus, whether a model like CLIP can achieve referential grounding through the pathway outlined in the argument from pre-training is doubtful.
+Consequently, CLIP lacks the selection history necessary to imbue its states with world-involving representational functions.
+It therefore fails to achieve referential grounding.
 By contrast with CLIP, some multimodal models are built with a pre-trained language model backbone.
-For example, typical VLMs that can receive images as well as text as input, such as Flamingo and LLaVA, leverage existing LLMs (with frozen weights) and learn a mapping between the visual encoder and the LLM (Bordes et al.
-2024).
+For example, typical VLMs that can receive images as well as text as input, such as Flamingo and LLaVA, leverage existing LLMs (with frozen weights) and learn a mapping between the visual encoder and the LLM (bordesIntroductionVisionLanguageModeling2024).
 This approach is computationally more efficient than training both vision and text encoders from scratch, as it only requires learning the connection between modalities.
-To the extent that the LLM backbone achieves referential grounding from pre-training on text alone, our framework entails that such model should inherit this property.
+To the extent that the LLM backbone achieves referential grounding from pre-training on text alone, our framework entails that such a model should inherit this property.
 The addition of a superficial mapping to the embedding space of a visual encoder provides a form of sensorimotor grounding, but it does not contribute to the referential grounding inherited from the pre-trained LLM.
 7.4 Embodiment
-Embodiment is often mentioned as the best way to secure that AI systems are grounded in the world, via their robotic bodies.
-Some embodied systems like SayCan (Ahn et al.
-2022) integrate a pre-trained LLM with a robotic body in a way that keeps these components functionally separate.
-In SayCan, an LLM breaks down verbal instructions into sequences of ’atomic’ behaviours that the robot can enact through low-level visuomotor control.
-Importantly, the LLM component is not further trained or fine-tuned based on the success or failure of the robot’s actions.
+Embodiment is often mentioned as the best way to ensure that AI systems are grounded in the world, via their robotic bodies.
+Some embodied systems like SayCan (ahnCanNotSay2022) integrate a pre-trained LLM with a robotic body in a way that keeps these components functionally separate.
+In SayCan, an LLM breaks down verbal instructions into sequences of ‘atomic’ behaviours that the robot can enact through low-level visuomotor control.
+Importantly, the LLM component is not further trained or fine-tuned based on the success or failure of the robot’s actions—there is no selection of internal states on the basis of their contributions to successful behaviour.
 In such systems, the embodied aspect adds nothing to the referential grounding of the language model component.
@@
 Like multimodality, physical embodiment alone does not automatically contribute to referential grounding.
-What matters is the causal-historical relationship established through training that selects for internal states with world-involving functional roles.
-Without such a learning history, embodiment remains external to the representational system itself.
 When fine-tuning is involved, the story is a bit different.
 Recent advances in robotic learning have led to the development of sophisticated multimodal systems that integrate vision, language, and action.
-Models such as RT-2 (Zitkovich et al.
-2023), RT-X (O’Neill et al.
-2024), and OpenVLA (Kim et al.
-2024) typically build upon language models or vision-language models that were pre-trained on text or multimodal data before being fine-tuned for robotic control.
+Models such as RT-2 (zitkovichRT2VisionLanguageActionModels2023a), RT-X (oneillOpenXEmbodimentRobotic2024), and OpenVLA (kimOpenVLAOpenSourceVisionLanguageAction2024) typically build upon language models or vision-language models that were pre-trained on text or multimodal data before being fine-tuned for robotic control.
 These models incorporate a pre-trained language model component and may inherit referential grounding from it.
-For instance, RT-2 builds upon vision-language models like PaLI-X and PaLM-E, which themselves incorporate language models pre-trained on text data.
-To the extent that these underlying language models achieve referential grounding through the mechanisms we have identified, the system as a whole may inherit this grounding.
+To the extent that the underlying language model achieves referential grounding through the mechanisms we have identified, the system as a whole may inherit this grounding.
 However, VLA models go beyond merely inheriting referential grounding from their language model components.
 During embodied training, they are specifically trained to map visual inputs and natural language instructions to robotic actions that achieve specific goals in the physical world.
-This additional training process plausibly establishes new functional roles for the model’s internal states that are directly tied to successful physical interactions.
-From our perspective, embodied training potentially enhances referential grounding by establishing or reinforcing causal-historical relations between internal states and features of the world that are relevant for successful manipulation.
-The training process selects for internal states that carry information about these features precisely because doing so improves performance on the physical tasks.
-This introduces an additional source of normativity tied to success in physical interaction, complementing whatever referential grounding was inherited from the pre-trained components.
+This additional training process may endow the model’s internal states with world-involving, representational functions grounded in a history of learning through physical interactions with the world.
 8 Concluding Remarks
-We have argued that the Symbol Grounding Problem, as introduced by Harnad (1990), can be adapted to apply to deep neural networks such as Generative Pre-Trained Models in we have called the Vector Grounding Problem.
-Despite initial intuitions, we have sought to demonstrate that at least a subset of current Generative Pre-Trained Models have the features necessary to be endowed with representations and outputs grounded in the world.
-While our discussion has focused on LLMs and, to a lesser extent, on their multimodal counterparts, any AI system fulfilling the causal-informational and historical relations requirements delineated above can possess representations with intrinsic meaning.
+We have argued that the Symbol Grounding Problem, as introduced by Harnad1990, also applies to deep neural networks such as generative pre-trained models, albeit in a slightly different form, which we have called the Vector Grounding Problem.
+We have sought to demonstrate that at least some current generative pre-trained models satisfy the requirements for having representations and outputs grounded in extra-linguistic referents.
+While our discussion has focused on LLMs and, to a lesser extent, on their multimodal counterparts, any AI system that fulfils the causal-informational and selection history requirements delineated above can possess representations with intrinsic meaning.
 Meeting these requirements is far from trivial, but nonetheless within reach of current AI models.
-As philosophical theories of representation have shown, in order to understand how cognitive systems – both biological and artificial – function, we must resist the temptation to assume that the presence of intrinsic meaning necessarily entails complex cognitive abilities such as understanding, knowledge, or consciousness.
-While having states imbued with intrinsic meaning might constitute one of the fundamental conditions for possessing these other cognitive abilities, it is insufficient on its own.
+As philosophical theories of representation have shown, to understand how cognitive systems—both biological and artificial—function, we must resist assuming that intrinsic meaning necessarily entails complex cognitive abilities such as understanding, knowledge, or consciousness.
+While having states imbued with intrinsic meaning might be a fundamental condition for possessing these cognitive abilities, it is insufficient on its own.
 Activations of specific populations of neurons in the mammalian cortex may represent the presence of edges in the visual scene, but these neuronal activations do not understand or possess knowledge about edges.
@@
 We argued for this claim in two steps.
-First, we distinguished between various forms of grounding, which are often conflated in the cognitive science and computer science literature.
-We suggested that one particular form of grounding, referential grounding, is central to both the Vector Grounding Problem and the older Symbol Grounding Problem.
-Referential grounding pertains to the way in which representations hook onto the world.
+First, we distinguished between various notions of grounding often conflated in the cognitive science and computer science literature.
+We suggested that one particular notion, referential grounding, is central to both the Vector Grounding Problem and the older Symbol Grounding Problem.
+Referential grounding pertains to how representations hook onto the world.
 It is the most fundamental type of grounding, and arguably serves as the basis for all other notions of grounding (with the partial exception of relational grounding).
-Referential grounding requires two main factors: causal-informational relations between internal states and the world, and historical relations that imbue those states with functions to carry information about the world.
+Referential grounding requires two main factors: causal-informational relations between internal states and the world, and a selection history that imbues those states with functions to carry information about the world.
 Second, we have shown that certain LLMs trained exclusively on linguistic data, particularly those that undergo preference tuning, are referentially grounded.
-These models benefit from indirect causal-informational relations with worldly entities and properties, mediated by human interactions with the world, which in turn shape the linguistic inputs they receive during training and inference.
+These models benefit from indirect causal-informational relations with worldly entities and properties, mediated by human interactions with the world that shape the linguistic inputs they receive during training and inference.
 Preference tuning introduces an additional learning objective that is directly world-involving, as it encompasses a requirement for truthfulness or factuality.
-The normativity of this requirement supplies an extra-linguistic accuracy standard for the representations and outputs of LLMs.
-More tentatively, we have suggested that even without preference tuning, the representations and outputs of LLMs may achieve referential grounding in some limited domains, through pre-training and in-context learning on specific kinds of input sequences, such as those describing game moves.
-This discussion led us to a counter-intuitive conclusion: some multimodal models trained on images in addition to text, like CLIP, are poor candidates for referential grounding, since they lack the kind of learning process that could supply world-involving functions.
-In fact, even some embodied systems that combine a ‘frozen’ LLM with a robotic body, like SayCan, also fail to acquire internal states and outputs that are referentially grounded.
-One benefit of our analysis is to highlight that multimodality and embodiment are neither necessary nor sufficient conditions for referential grounding in artificial systems, although they are certainly compatible with it if the right kind of learning process is available.
+The normativity of this requirement supplies an extra-linguistic accuracy standard that selects and stabilises the internal states that contribute to meeting it.
+More tentatively, we suggested that even without preference tuning, the representations and outputs of LLMs may achieve referential grounding in limited domains through pre-training and in-context learning on specific input sequences, such as those describing game moves.
+This discussion led to a counter-intuitive conclusion: some multimodal models trained on images in addition to text, like CLIP, are poor candidates for referential grounding, since they lack the kind of learning process that could supply world-involving functions.
+Even embodied systems that combine a ‘frozen’ LLM with a robotic body, like SayCan, fail to acquire internal states and outputs that are referentially grounded.
+Our analysis highlights that multimodality and embodiment are neither necessary nor sufficient conditions for referential grounding in artificial systems, although they are certainly compatible with it given the right kind of learning process.